### v3

# Towards Generalised Pre-Training of Graph Models

Abstract
The principal benefit of unsupervised representation learning is that a pre-trained model can be fine-tuned where data or labels are scarce. Existing approaches for graph representation learning are domain specific, maintaining consistent node and edge features across the pre-training and target datasets. This has precluded transfer to multiple domains.
In this work we present Topology Only Pre-Training, a graph pre-training method based on node and edge feature exclusion. Separating graph learning into two stages, topology and features, we use contrastive learning to pre-train models over multiple domains. These models show positive transfer on evaluation datasets from multiple domains, including domains not present in pre-training data. On 75% of experiments, ToP models perform significantly ($P\leq 0.01$) better than a supervised baseline. These results include when node and edge features are used in evaluation, where performance is significantly better on 85.7% of tasks compared to single-domain or non-pre-trained models. We further show that out-of-domain topologies can produce more useful pre-training than in-domain. We show better transfer from non-molecule pre-training, compared to molecule pre-training, on 79% of molecular benchmarks.
1 Introduction
Much work has been conducted on forming useful representations of graphs which are more easily used by well-explored methods. Khoshraftar et al. [15], Hamilton et al. [12] and Chen et al. [4] survey graph representation learning in general. Such representation learning, after transfer and fine-tuning, allows greater performance in areas of data or label scarcity.
Whilst there is a good deal of research into graph representation learning methods, there is little work on methodologies to combine different domains of graph data effectively. Ubiquitous in other fields, primarily text and images, is pre-training on a large amalgamated dataset of samples, aiming to span the full or realistic space of the data [28]. The main obstacle to such an approach for graphs is consistency in node and edge features, as graph data presents a large diversity in possible feature-sets.
Given this variety of features between domains current research argues, but does not show, that multi-domain pre-training on graphs is not feasible [21, 31, 39]. There are very few works attempting learning across multiple domains of graphs [19]. This excludes areas of data or label-sparsity from the benefits of pre-training. Multi-domain pre-training will allow learning on domains where data is too scarse to form pre-training datasets.
Here we propose Topology Only Pre-Training (ToP), a generalised pre-training method for graph models, under which pre-training is conducted with node and edge features excluded. We create a multi-domain dataset, and using ToP through adversarial contrastive learning, produce pre-trained models with consistently positive transfer across multiple domains and tasks. Transfer is also positive when node and edge features are included in downstream datasets during fine-tuning.
In this work we propose and investigate ToP as a pre-training method for graph data that is not limited to a given domain or task. The result of such pre-training is a single model that can be transferred with broadly positive performance increases on arbitrary graph tasks and domains. Experimentally we test two hypotheses:
Hypothesis 1: Excluding Features
Pre-training on multiple domains with node and edge features excluded can lead to consistent positive transfer compared to fully-supervised, task-specific models.
Hypothesis 2: Sample Diversity
With the exclusion of node and edge features out-of-domain pre-training can have greater benefits than in-domain pre-training across task-relevant metrics.
Findings We evaluate ToP using graph constrastive learning with both random and adversarial augmentations. Through this evaluation we contribute the following empirical findings:
  1. 1.
    Effective pre-training without features. Pre-training on varied domains, under the exclusion of node and edge features, leads to positive transfer. We find that under fine-tuning, results are at worst on-par, and on 76% of tasks significantly better, than a non-pretrained baseline.
  2. 2.
    Increased performance from domain diversity. Pre-training without features from varied non-target domains, instead of the target domain, can bring greater downstream performance benefits. On molecular benchmarks, we find that a model pre-trained on non-molecules, compared to pre-trained only on molecules, is at worst on-par. On 56% of tasks the non-molecule model performs significantly better, and never significantly worse, than the molecule model.
Given these findings, under recent definitions, these may be among the first foundation graph models [2, 19, 21, 39]. Both of these findings go beyond the current consensus of graph pre-training, opening broad avenues of research and application. In particular our second finding runs contrary to the single-domain training paradigm.
Organisation In Section 2, we describe the notation and concepts necessary to understand our method and provide an overview of related works. In Section 3, we describe our generalised pre-training method ToP. In Section 4, we describe how we evaluate the hypotheses detailed above, then detail the experiments we perform to evaluate ToP and these hypotheses. In Section 5 we discuss the implications of the results from these experiments, as well as list the assumptions we make and threats to validity. In Section 6, we conclude with avenues for future research.
2 Background and Related Works
Here we describe the concepts necessary to understand the rest of this work.
Representation Learning Let $X\in\mathcal{X}$ represent a random variable $X$ in the data space $\mathcal{X}$, and $Y\in\mathcal{Y}$ the same for the output space. Representation learning aims, given their joint distribution $P(X,Y)$, to learn a mapping $X\rightarrow\hat{X}$ that maximises their mutual information $I(X;Y)$. In most contexts this means producing numerical vectors for input data that usefully describe that data.
Contrastive Learning Let $\mathcal{T}$ be set of small transformations called ‘augmentations’, where $t\in\mathcal{T}:X\rightarrow X^{\prime}$. These augmentations are generally designed to manipulate the input sample in a way non-destructive to the ‘useful’ information it represents. Common choices for images, for example, are gray-scaling and rotations. Given an encoder’s mapping $F(X_{i})\rightarrow\hat{X_{i}}$, multiple augmentations of the same input sample should be as similar as possible.
Graph Representation Learning Hamilton [11] provides a comprehensive summary of methods for Graph Representation Learning (GRL). Traditional methods to encode graphs as vectors include graph statistics such as centrality measures and clustering coefficients, but also kernel methods such as the Weisfeiler-Lehman (WL) kernel [29]. Expressive Graph Neural Networks (GNNs) models, generally through parameterised message-passing, can produce powerful representations of graphs. A number of GNN architectures have been proposed with the Graph Isomorphism Network (GIN) being the most expressive [41] in representation learning contexts.
Graph Contrastive Learning Multiple studies exist for graph contrastive learning (GCL) [5, 10, 13, 14, 18, 38, 42]. The typical graph contrastive learning formulation, Graph Contrastive Learning (GraphCL), is put forward by You et al. [38]. In GraphCL and many subsequent works, such augmentations are performed randomly using fixed parameters, i.e., ‘drop 20% of nodes’.
Adversarial Graph Contrastive Learning It has been shown that the quality of augmentations is critical for graph contrastive learning [38, 42]. To maximise expressivity in these augmentations adversarial strategies have been proposed [13, 40]. Suresh et al. [32] propose Adversarial-Graph Contrastive Learning (AD-GCL), with their method training both an encoder and a ‘view-learner’ $\mathcal{T}_{\theta}$. AD-GCL performs the standard GCL protocol with weight updates to a parameterised ($\theta$) set of augmenting functions $\mathcal{T_{\theta}}$ through the contrastive loss. Here $\mathcal{T_{\theta}}$ learns augmentations that minimise mutual information between the source and augmented graphs through edge dropping. The encoder, adversarially to this, learns to maximise shared information between the respective augmentations. The updates for $\mathcal{T}$ are regularised such that augmentations are not entirely destructive.
One For All A work concurrent with ours is One-For-All (OFA) [19]. OFA employs a single model on graph and node classification tasks across domains, sidestepping the issues around node and edge features we’ve discussed by assuming that natural language descriptions of features are adequate, and using language model embeddings as features. OFA, then, requires that input graphs have textual features, or that existing features can be usefully described in natural language. Whereas OFA produces only classification models, ToP produces pre-trained models intended for downstream fine-tuning. As such ToP allows arbitrary downstream tasks to be conducted which are not possible with OFA, regression being an example.
3 ToP: Topology Only Pre-Training
Figure 1: The process for handling features with ToP. During pre-training, features are replaced with a single integer label for each node and edge (all identical), which are then passed to a single-layer MLP. This results in an identical input vector for each node. During transfer, a different input head is used, moving the original dimensions to the hidden dimensionality of the GNN encoder block. This allows arbitrary node and edge features to be included in transfer. Omitting the output head of the model allows transfer onto node and edge-level tasks.
One naive approach to generalised pre-training with varied feature sets included would be to pass features as-is, with padded vectors where feature sets are of a lower dimensionality. Here components would be shared across domain features, for example types of atoms and number of social followers in the same component, representing an extremely difficult space to learn. Alternatively one could reserve components of a long feature vector for each feature set, with padding tokens in components for other domains and feature sets. This would require a model to learn a massive span of features, and preclude movement to new feature sets, and again learning is unlikely to occur. Feature-management approaches, which move domain features to the same feature set [19], are in their infancy.
Given that pre-training while including features in their original form is not feasible, we view graph learning as a task occurring on both topologies $G$ and features on that topology $X$. From this view, pre-training without features is well-motivated, with models learning generalised information about graph structures before features are introduced. This information from $G$ is useful for downstream tasks, even while missing task-specific information, and in situations where data is abundant pre-training on both $X$ and $G$ will likely still lead to higher performance.
This is an abstraction away from feature-specific correlations, with abstraction a core quality of effective representation learners [16]. Pre-training without topologies, viewed as an initialisation, would provide a smoother loss landscape for downstream tasks [20]. In this case a wider range of topologies could represent more useful information than a narrow range closer to a downstream domain.
Topology Only Pre-Training (ToP) is a method for generalised pre-training of graph models. Under ToP pre-training is conducted only on topologies, with node and edge features excluded or replaced with identical values. This use of only topologies $G$, instead of graphs with specific features $\underline{G}$, allows pre-training on a multi-domain dataset. Prior to this work, the use of such a graph dataset with mixed node and edge features was not considered feasible [21, 31, 39].
3.1 Hypothesis 1: Excluding Features
Supporting Hypothesis 1 requires producing a model that has learnt useful representations of graphs in general. Our first step to achieving this is to construct a dataset that covers a variety of graph domains. Such a dataset should contain graphs that present a large diversity of multi-node patterns. Our approach is simple: include graphs from several domains and assume that there is adequate coverage of the multi-node patterns that occur in real-world graphs for a given set of downstream tasks. The same approach has been taken in the domains of text [3] and images [24], where enormous numbers of data samples are scraped from the web for generative pre-training. This relies on two assumptions 1) this dataset adequately covers the relevant space for downstream transfer, and 2) a single model can usefully represent such a space. These assumptions, as in other fields, are justified through performance metrics on downstream tasks on seen and unseen domains.
We formulate our test for Hypothesis 1, that pre-training on graphs without node or edge features is beneficial, as beating a simple supervised baseline. Said baseline is a model with the same architecture as the pre-trained model but without pre-training. Both models are then trained, or fine-tuned for the pre-trained model, on the same downstream datasets under the same conditions. On each downstream task, should the pre-trained model match or out-perform the supervised model, this represents positive transfer. Positive transfer on a task indicates that at least some of the learning during pre-training was beneficial to subsequent learning during fine-tuning. A positive transfer rate across datasets of more than 50% indicates that in-general transfer is positive, thus meeting the minimum criteria for Hypothesis 1. A high positive transfer rate strongly supports Hypothesis 1.
3.2 Hypothesis 2: Sample Diversity
From Hypothesis 2 we expect that for a given task models trained on a more diverse pre-training domain can bring better transfer than a model trained on the target domain. We do not argue that this is always true, only that in some settings out-of-domain samples can be more useful in pre-training than in-domain. This has clear parallels to the performance-by-scale approach taken in other fields [28, 6, 23]. To this end we compare two models, the first trained on a dataset from a single domain, as is the current practice. This single domain matches that of the downstream target dataset(s). The second model is instead trained on the rest of the data while excluding the target domain.
The total size of the pre-training datasets for each model should match for fair comparison. By pre-training both models under the same controlled conditions (hyperparameters etc.), then evaluating their performance under fine-tuning, we directly compare pre-training on a single domain to pre-training on multiple domains. General positive transfer from the multi-domain model, in comparison to transfer from the in-domain model, would constitute a clear supporting evidence for Hypothesis 2.
4 Experiments
Our experiments are guided by two hypotheses concerning ToP.
4.1 Datasets
We begin by collating a collection of graph datasets. Statistics are available in the Supplemental Material in Table 3. Here we provide qualitative detail as well as any downstream task information:
molpcba [14, 35]
A curated dataset of 437,929 small molecules, from which we use 250,000. Used only in pre-training.
Cora [22]
A much used citation graph. Used only in pre-training.
Fly Brain [34]
The full neural connectome of a fruit fly larvae. As the multi-graph is dense at originally 548,000 edges, we include a single edge between neurons if there are more than two synapses between them. Used only in pre-training.
Facebook [25]
A single graph of page-page connections on Facebook. Used both in pre-training and downstream transfer. On a graph level the downstream task is predicting average clustering, and we also evaluate transfer to node classification.
Twitch [26]
Ego networks from the streaming platform Twitch, used both in pre-training and downstream transfer, with a downstream binary classification task.
Roads [17]
The road network of Pennsylvania. Used both in pre-training and downstream transfer, with the task set to be predicting graph diameter normalised by graph size.
Benchmark Molecules [14]
molesol, molfreesolv, mollipo, molclintox, molbbbp, molbace, molhiv. Smaller datasets sampled from MoleculeNet, with single-target classification or regression downstream tasks. These are used only in downstream transfer.
Trees
Trees, of a randomly sampled maximum depth, and a randomly fixed probability of branching at each level, with a downstream task of predicting depth. Used only in downstream evaluation.
Random
Erdos-Renyi (ER) graphs [7], with a number of nodes and edge probability randomly sampled, and predicting edge probabilities as a downstream task. Used only in downstream evaluation.
Community
Small ER subgraphs, with a set intra-subgraph and random inter-subgraph connection probability, with inter-subgraph connectivity as target. Used only in downstream evaluation.
As the Facebook, Cora, Roads and Fly Brain datasets consist of a single graph, we employ exploration sampling to construct datasets of many smaller graphs. We use a randomly selected exploration sampler [^1] Using the littleballoffur Python library [27]. for each sample, which in turn samples a random number of nodes in the range $24\leq|V|\leq 96$ from the source graph. The use of a selection of exploration samplers should further ensure that a large variety of graphs and topological features are included in the datasets.
Figure 2: A UMAP embedding of encodings from each model, as well as an untrained model. The untrained and Chem models are noticeably more fragmented than the Social and All models. In turn the Social model is more fragmented than the All model. In the All model embedding, molecules (primarily dark blue) are all in the same region, furthest from the Twitch Ego networks (lilac), with the Cora (brown) and Facebook Page-Page data (pink) “filling the gap”.
We take the downstream task on each validation set where available and add their relevant scores (MSE and $1-$ AUROC) with a linear model as a crude monitor for the models’ representational abilities. We conduct a limited Bayesian hyperparameter sweep. The heuristic used is addition of evaluation scores on each dataset weighted by the number of samples in that dataset.
4.2 Models
We perform a limited ablation study by evaluating pre-training five models on different subsets of the amalagamated dataset: 1) GIN, an untrained GIN model of identical architecture. 2) ToP-Chem, an AD-GCL model fit only on the molecular molpcba dataset. 3) ToP-Social, an AD-GCL model fit only on non-molecular datasets (Cora, Facebook, Fly Brain, Roads, Twitch). 4) ToP-All, an AD-GCL model fit on all the training data. 5) All (CL), a model trained with the loss function from You et al. [38] using random edge dropping ($\lambda=0.2$).
Models during pre-training, unless specified, are six GIN [37] layers with single linear layers as input and output. The GIN model acts as our supervised baseline (Hypothesis 1). The ToP-Chem/ToP-Social models indicate the impact of training data diversity on downstream transfer performance, particularly on our downstream molecular datasets (Hypothesis 2). The ToP-All (CL) model shows the benefits of adversarial augmentations.
Following the results of the hyperparameter sweep, we train each model for 100 epochs with a batch size of 512, learning rates of 0.001, embedding and projection dimensions of 300 and a regularisation constant of 0.2. Figure 2 shows a UMAP embedding of encodings of the validation datasets for an untrained model and the pre-trained models. A larger version can be found in the Supplemental Material.
Pre-training was conducted on a workstation with an NVIDIA A6000, using around 16GB of video memory capacity. Pre-training the ToP-All model took a little more than eight hours. Transfer results were produced on consumer laptops, primarily with an Apple M1 processor, and on GPU use around 3GB of video memory with batch sizes of 128.
4.3 Transfer
Here we apply our pre-trained models to various downstream tasks through transfer learning. This follows stages of abstraction away from ToP pre-training: first we fine tune on datasets without features, then on datasets with features included, and finally on node and edge-level tasks.
Table 1: Scores for fine-tuning on each validation dataset. For the classification datasets we report AUROC and for regression datasets RMSE. Bold text indicates a result beats the unsupervised (GIN) baseline. Underlined text indicates a superior result, down to three significant figures. Checkmarks ✓indicate that the ToP-All model significantly out-performs the GIN baseline at $p<0.01$. ‘Best’ indicates the proportion of datasets over which a model has superior performance, and ‘Positive’ the proportion of datasets over which performance is superior to the fully-supervised baseline.
GIN (Baseline) All (CL) ToP-Chem ToP-Social ToP-All
molfreesolv 4.98 $\pm$ 0.48 4.20 $\pm$ 0.10 3.89 $\pm$ 0.21 3.83 $\pm$ 0.11 4.26 $\pm$ 0.03
molesol 1.78 $\pm$ 0.09 1.57 $\pm$ 0.06 1.44 $\pm$ 0.09 1.47 $\pm$ 0.04 1.31 $\pm$ 0.03
mollipo 1.13 $\pm$ 0.04 1.07$\pm$ 0.03 1.04 $\pm$ 0.02 0.99 $\pm$ 0.01 0.97 $\pm$ 0.01
molclintox 0.53 $\pm$ 0.04 0.46 $\pm$ 0.03 0.51 $\pm$ 0.03 0.50 $\pm$ 0.05 0.55 $\pm$ 0.11
molbbbp 0.57 $\pm$ 0.04 0.60 $\pm$ 0.03 0.54 $\pm$ 0.04 0.60 $\pm$ 0.03 0.59 $\pm$ 0.02
molbace 0.68 $\pm$ 0.05 0.76 $\pm$ 0.03 0.63 $\pm$ 0.13 0.75 $\pm$ 0.03 0.71 $\pm$ 0.02
molhiv 0.36 $\pm$ 0.06 0.50 $\pm$ 0.14 0.54 $\pm$ 0.13 0.64 $\pm$ 0.05 0.66 $\pm$ 0.04
Facebook 0.31 $\pm$ 0.07 0.24 $\pm$ 0.06 1.17 $\pm$ 0.17 0.13 $\pm$ 0.00 0.13 $\pm$ 0.00
Twitch Egos 0.72 $\pm$ 0.01 0.74 $\pm$ 0.01 0.69 $\pm$ 0.02 0.74 $\pm$ 0.01 0.79 $\pm$ 0.01
Roads 0.46 $\pm$ 0.07 0.25 $\pm$ 0.03 0.57 $\pm$ 0.07 0.10 $\pm$ 0.00 0.10 $\pm$ 0.00
Trees 0.20 $\pm$ 0.01 0.19 $\pm$ 0.02 0.24 $\pm$ 0.03 0.12 $\pm$ 0.00 0.12 $\pm$ 0.00
Community 0.74 $\pm$ 0.11 0.31 $\pm$ 0.07 0.72 $\pm$ 0.13 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00
Random 0.53 $\pm$ 0.23 0.29 $\pm$ 0.10 0.59 $\pm$ 0.09 0.04 $\pm$ 0.00 0.04 $\pm$ 0.00
Best 0% 15.4% 0.0% 38.5% 69.2%
Positive 92.3% 38.5% 92.3% 100%
Features Excluded We replace the output head with a simple two-layer MLP, with a single output node. We then perform 10 fine-tuning runs on each validation dataset, with and without features, for each pre-trained model. Table 1 presents results of transfer and fine-tuning without features.
The ToP-All model consistently out-performs the other models under fine-tuning on the majority of datasets. At $p\leq 0.01$ the ToP-All model significantly out-performs the GIN baseline on 77% of datasets. When performance is not significantly better, it is never significantly worse. Notably the ToP-All model always out-performs the All (CL) model, indicating the merits of using a learnt augmentation scheme in this multi-domain setting. The standard deviation of the results on each dataset are lower than those for the non-pretrained model and the domain-specific pre-trained models. In Figure LABEL:fig:training-examples in the supplemental material we show validation loss on the several datasets during fine-tuning for the ToP-All model and an un-pre-trained GIN with the same architecture. Here the advantages of pre-training are clear: the pre-trained model has consistently lower validation loss, seems to plateau later and at a lower loss, and has a lower deviation in that loss than the baseline GIN.
Pre-training only on molecules (‘Chem’ data subsets) results at-best on-par transfer compared to the supervised GIN baseline, with negative transfer on the majority of datasets. Pre-training only on non-molecules, in comparison, results in positive transfer on the large majority of datasets and pre-training methods. At $p\leq 0.01$ the ToP-Social model significantly out-performs the ToP-Chem model on the majority of datasets. As before, where performance is not significantly better, it is never significantly worse. We employ an ablation study over the non-molecular pre-training datasets (see Supplemental Material) to ensure that this is not a single artifact result. This study corroborates the results presented in Table 1, with the Chem model consistently out-performed by non-molecular models.
We find that under pre-training with random edge dropping performance actually drops when including molecular data (‘Social’ vs ‘All’). This suggests that AD-GCL is able to differentiate between molecules and non-molecules and applies different augmentations to these domains. Nonetheless, we show that using a variety of non-molecular graphs in-general provides a more solid pre-training foundation than including molecules for both methods of augmentation.
Table 2: Scores for fine-tuning pre-trained GIN models (and a non-pre-trained baseline) on the molecular evaluation datasets with node and edge features included via projection heads. Results are reported in the same manner as in Table 1.
GIN (Baseline) All (CL) ToP-Chem ToP-Social ToP-All
molfreesolv 4.00 $\pm$ 0.43 4.00 $\pm$ 0.24 4.16 $\pm$ 0.29 3.61 $\pm$ 0.43 3.01 $\pm$ 0.59
molesol 1.72 $\pm$ 0.25 1.18 $\pm$ 0.09 1.62 $\pm$ 0.31 1.03 $\pm$ 0.17 0.90 $\pm$ 0.04
mollipo 1.09 $\pm$ 0.05 0.99 $\pm$ 0.05 1.09 $\pm$ 0.06 0.80 $\pm$ 0.02 0.79 $\pm$ 0.02
molclintox 0.65 $\pm$ 0.12 0.56 $\pm$ 0.16 0.52 $\pm$ 0.03 0.86 $\pm$ 0.02 0.84 $\pm$ 0.05
molbbbp 0.60 $\pm$ 0.03 0.63 $\pm$ 0.04 0.61 $\pm$ 0.03 0.62 $\pm$ 0.03 0.67 $\pm$ 0.02
molbace 0.75 $\pm$ 0.03 0.72 $\pm$ 0.05 0.70 $\pm$ 0.07 0.70 $\pm$ 0.05 0.72 $\pm$ 0.05
molhiv 0.32 $\pm$ 0.03 0.46 $\pm$ 0.14 0.37 $\pm$ 0.07 0.62 $\pm$ 0.11 0.57 $\pm$ 0.13
Best 14.3% 0.0% 0.0% 28.6% 57.1%
Positive 57.1% 42.9% 85.7% 85.7%
Features Included By swapping the input head of the pre-trained model, as seen in Figure 1, arbitrary node and edge features can be included during transfer. To this end we replace the input head of our model with a three layer Multi-Layer Perceptron (MLP) for both node and edge features, then perform fine-tuning runs on our molecular datasets. Table 2 presents the results of this transfer and fine-tuning. These tasks are assumed to require node and edge features to obtain good performance, and here positive transfer rates are similar to when no features are included, despite ToP’s exclusion of features during pre-training. The ToP-All and ToP-Social models again demonstrate consistent positive transfer, and on most datasets their performance improves compared to when features were excluded. The same results are presented for different GNN backbones in Table 9 in the Supplemental Material, but perform no optimisation for these models, using hyperparameters from the GIN models here for the GAT and GCN models presented. As such these results are included only for completeness.
Node & Edge Transfer By excluding the output head (again see Figure 1) we are able to transfer ToP pre-trained models onto node and edge level tasks. Here we focus on edge prediction and node classification, and though ToP pre-training does seem to offer some performance increase, the benefits are far smaller than on graph level tasks. These results can be found in the Supplemental Material in Table 10.
Further results can be found in the Supplemental Material in Section B: see Table 5 for transfer with a linear models, Tables (6,7) for GraphCL-like augmentations, Table 8 for a limited ablation over non-molecular pre-training datasets, Table 9 for ToP with varied GNN backbones, and Table 10 for node and edge classification. A comparison of ToP against domain-specific methods is in Table 11.
5 Discussion
Here we discuss our findings in terms of our hypotheses. This includes discussion on the performance of, and reasonable use cases for, models trained with ToP. Finally we detail explicitly assumptions we make and potential risks to the validity of this work.
5.1 Findings
Pre-training on multiple domains of topologies offers significant performance improvements on downstream tasks (Hypothesis 1). With features excluded we find entirely positive transfer across a range of domains and graph-level tasks using adversarially pre-trained models with ToP. With features included, on popular molecular benchmarks, there are significant performance increases on graph-level tasks compared to fully supervised models.
Further, exposure to a larger variety out-of-domain samples can benefit performance. In Table 1, assuming that the ToP-Chem model has the least varied training data, then the ToP-Social model, then the ToP-All model, performance increases along the same order (Hypothesis 2). Performance between the ToP-Social and ToP-All models is highly comparable, despite the ToP-All model having double the training samples. We expect, though we leave this as an area for future research, that a yet greater breadth of domains in the pre-training data would bring greater performance benefits.
The most intuitive assumption would be that exposing a molecular representation learner to non-molecular graphs would hinder its performance. Instead the ToP-Social and ToP-All models out-perform the ToP-Chem model almost all molecular datasets under fine-tuning. This holds true for varied pre-training regimes and model backbones (see Tables 67 and 9 in the Supplemental Material). This shows potential for future work using ToP. Pre-training on graphs could be possible through data scale and diversity in-general, instead of within a specific domain. This has clear parallels to the current trends in generalised pre-training in other fields [23, 33].
Similarly one might expect that including node and edge features during fine-tuning (or supervised training of a non-pretrained model) would likely lead to these domain-specific models out-performing all-domain models. Our results instead show that performance is both superior and more consistent from our all-domain model on featured molecular graphs, with performance increasing compared to the same model on non-featured graphs. This allows models, after using ToP, to be transferred to arbitrary tasks on graph structures, without the inherent costs of LLM-based approaches [19].
Given these findings the reasonable use-cases for models pre-trained with ToP are broad. The potential benefit is two-fold. Firstly, pre-training and fine-tuning requires less data for the downstream task to be viable. This is particularly valuable in areas of data scarcity or expense, as well as offering presumed decreases in training time for the final model. Secondly, as shown by our results in this work, our pre-trained models offer actual performance increase over simple supervised training. Our assumption is that the broad range of generic topological features learnt by the ToP model provide a more robust foundation for training than simple random initialisation, so pre-trained models avoid overfitting to non-useful features better than randomly initialised models. Guarantees on this point are left for future research.
5.2 Assumptions, Risks, Future Directions
The most significant assumption we made — that a model can learn representations that are useful on multiple domains of downstream tasks — has been justified by our results. However, in the full scope of downstream tasks for graph data, this first exploratory work is still fairly limited. Firstly we do not experiment with novel graph augmentations (or “views”), instead employing only node and edge dropping. While we’ve shown that adversarial augmentations do improve pre-training, adversarial augmentations could be developed that include some techniques present in Hassani and Khasahmadi [13], for example edge addition alongside edge dropping. Recent works have found that such augmentations, including combined additive and subtractive edge-based augmentations, can lead to more robust representation learners [9]. We anticipate that a variety of augmentations, paired with a more detailed study of how different compositions of domain data influence representations, could bolster the expressivity of these topology-pre-trained models.
Secondly our experiments are limited to a few set compositions of domains. The non-task-specific evaluation of representation models is a challenging and open research problem. The same challenge is compounded on the high-dimensional graph space, which in comparison to other data structures both lacks intuitive human understanding and has a smaller body of existing research. The usual workaround for this issue is increasing the scale and breadth of the pre-training data [28]. Future work using ToP is likely to follow this same path.
Additionally this work primarily focusses on graph-level tasks, in the same manner as the original AD-GCL work [32]. We anticipate that future work will optimise ToP, probably through targeted augmentation schemes, for node and edge related tasks specifically.
Lastly we should note that we haven’t compared ToP to other graph pre-training methods. Such methods are, with a few exceptions [19], domain specific. As such comparisons would be between our generalised models and models pre-trained on a given domain with all features included during pre-training. Such pre-training requires a large dataset of labelled samples, significantly limiting viable domains. The Supplemental Material contains a comparison of ToP pre-trained models with AD-GCL [32], GraphCL [13] and InfoGraph [36].
Liu et al. [21] recently proposed definitions for, and expected properties of, graph foundation models. The models produced with ToP satisfy their broad definition, ‘…benefit from the pre-training of broad graph data, and can be adapted to a wide range of downstream graph tasks’. The two further properties proposed by Liu et al. [21], emergence and homogenization, are less easy to demonstrate. Liu et al. themselves note that how these properties manifest is an open question, and therefore we similarly leave whether ToP models exhibit these properties as an area for future research.
6 Conclusion
We presented ToP, a method for generalised multi-domain pre-training of graph models. ToP relies on excluding node and edge features during pre-training. Features can then be re-included during fine-tuning on downstream tasks.
Findings Through contrastive learning we used ToP to present graph models capable of positive transfer across domains and tasks. These are an important step towards graph foundation models with use-cases in the manner of models such as BERT [6]. Like BERT, an early foundation language model, ToP models can be fine-tuned with positive transfer on multiple domains and tasks. On 75% of downstream tasks from multiple domains we show significant ($p\leq 0.01$) positive transfer compared to a non-pre-trained model with the same architecture, as expected from our first hypothesis. This is also true when node and edge features are re-introduced during fine-tuning. Compared to a supervised baseline, where results are not significantly better, they are never significantly worse.
Additionally we showed that ToP pre-training on a single target domain leads to at-best on-par, and on most tasks significantly worse ($p\leq 0.01$), performance compared to pre-training on multiple non-target domains. This clearly demonstrates that the variety provided by out-of-domain pre-training can hold more value than the status-quo in-domain pre-training, as in our second hypothesis.
Directions We anticipate that future work will focus on several key areas: 1) Improved or tailored graph augmentation strategies. 2) Investigation of domain weighting in dataset compositions. 3) A larger pre-training dataset spanning a wider range of domains. 4) Improved performance on node-level and edge-level tasks.
Appendix A Supplemental Material
A.1 Dataset Statistics
Table 3: Statistics for each of our datasets. If used for training, we denote these statistics averaged across the training data. Otherwise we report statistics for their validation (downstream fine-tuning) datasets. ‘BC’ indicates binary classification.
Num. Graphs Number of Nodes Number of Edges Diameter Avg. Clustering Downstream Task
Training Only
molpcba 250000 25.7 $\pm$ 6.28 27.7 $\pm$ 7 13.5 $\pm$ 3.29 0.00128 $\pm$ 0.0115
Cora 50000 59.3 $\pm$ 20.7 119$\pm$ 56.6 10.1 $\pm$ 4.68 0.322 $\pm$ 0.0845
Fly Brain 50000 59.5 $\pm$ 20.8 146$\pm$ 99.6 8.8 $\pm$ 3.49 0.237 $\pm$ 0.0875
Training and Evaluation
Twitch 50000 29.6 $\pm$ 11.1 86.6 $\pm$ 70.7 2$\pm$ 0 0.549 $\pm$ 0.149 BC. Games Played
Facebook 50000 59.5 $\pm$ 20.7 206 $\pm$ 170 10.2 $\pm$ 6.39 0.429 $\pm$ 0.133 Avg. Clustering
Roads 50000 59.5 $\pm$ 20.8 73.9 $\pm$ 28.2 16.4 $\pm$ 5.92 0.0559 $\pm$ 0.0426 Diameter / Num. Nodes
Evaluation Only
molesol 1015 12.7 $\pm$ 6.64 12.9 $\pm$ 7.62 6.81 $\pm$ 3.30 0.000353 $\pm$ 0.00417 Water Solubility
molfreesolv 577 8.46 $\pm$ 3.97 8.02 $\pm$ 4.50 5.02 $\pm$ 2.09 0.0 $\pm$ 0.0 Solvation
mollipo 3780 27.0 $\pm$ 7.44 29.4 $\pm$ 8.22 13.8 $\pm$ 4.03 0.00366 $\pm$ 0.017 Lipophilicity
molclintox 1329 26.3 $\pm$ 15.8 28.0 $\pm$ 17.1 12.4 $\pm$ 6.09 0.00259 $\pm$ 0.0191 BC. Toxicity
molbbbp 1835 23.5 $\pm$ 9.89 25.4 $\pm$ 11.0 11.2 $\pm$ 4.03 0.00284 $\pm$ 0.0278 BC. Penetration
molbace 1361 34.0 $\pm$ 7.89 36.8 $\pm$ 8.13 15.2 $\pm$ 3.14 0.00672 $\pm$ 0.0205 BC. Bioactivity
molhiv 37014 25.3 $\pm$ 12.0 27.3 $\pm$ 13.1 12.0 $\pm$ 5.15 0.00158 $\pm$ 0.0156 BC. Inhibition
Evaluation Only, Unseen Domain
Trees 5000 19.6 $\pm$ 6.96 18.6 $\pm$ 6.96 10.1 $\pm$ 3.11 0$\pm$ 0 Depth
Random 5000 29.3 $\pm$ 10.7 111$\pm$ 82.1 3.88 $\pm$ 1.16 0.227 $\pm$ 0.0998 Density
Community 5000 48$\pm$ 0 323$\pm$ 26.2 3$\pm$ 0.0346 0.408 $\pm$ 0.0256 Inter-Community Density
A.2 Exploration of Embedding Space
Figure 3: A UMAP embedding of encodings from each model, as well as an untrained model. The untrained and Chem models are noticeably more fragmented than the Social and All models. In turn the Social model is more fragmented than the All model. In the All model embedding, molecules (primarily dark blue) are all in the same region, furthest from the Twitch Ego networks (lilac), with the Cora (brown) and Facebook Page-Page data (pink) “filling the gap”.
On the combined encodings of the validation sets we compute $R^{2}$ correlation coefficients between the first five principal components (PCA) and common graph-level metrics. High correlations, or the lack thereof, should give some indication of how embedding components represent different graph characteristics. The results are presented, along with each component’s explained variance, in Table 4 and visualised in Figure 4.
Table 4: PCA components, fit on the embeddings of the test set, and their most correlated graph-level metric, for each of the trained models. Underlined text indicates a strong correlation $|R^{2}|\geq 0.66$, and bold text indicates moderate correlation $0.33\leq|R^{2}|\leq 0.66$. We show up to the first five PCA components, but include only three where explained variance values are minimal.
Variance Ratio Metric Correlations
Most $R^{2}$ Second $R^{2}$ Third $R^{2}$
Untrained
PCA 0 $9.96\times 10^{-01}$ Edges $0.219$ Transitivity $0.072$ Density $0.063$
PCA 1 $3.52\times 10^{-03}$ Edges $-0.555$ Transitivity $-0.276$ Density $-0.243$
PCA 2 $1.13\times 10^{-04}$ Edges $0.39$ Transitivity $0.224$ Density $0.21$
Chem
PCA 0 $9.97\times 10^{-01}$ Edges $0.223$ Transitivity $0.074$ Density $0.065$
PCA 1 $2.60\times 10^{-03}$ Edges $-0.514$ Transitivity $-0.253$ Density $-0.225$
PCA 2 $1.14\times 10^{-05}$ Edges $0.132$ Nodes $0.051$ Transitivity $0.047$
Social
PCA 0 $9.99\times 10^{-01}$ Edges $0.127$ Transitivity $0.032$ Density $0.032$
PCA 1 $9.92\times 10^{-04}$ Nodes 1.00 Edges $0.528$ Density $-0.393$
PCA 2 $4.79\times 10^{-08}$ Edges $0.133$ Density $0.07$ Transitivity $0.054$
All
PCA 0 $7.04\times 10^{-01}$ Nodes 0.999 Edges $0.534$ Density $-0.388$
PCA 1 $1.94\times 10^{-01}$ Edges 0.805 Density 0.737 Diameter $-0.554$
PCA 2 $4.23\times 10^{-02}$ Clustering $0.309$ Transitivity $0.248$ Edges $0.235$
PCA 3 $2.43\times 10^{-02}$ Diameter $0.5$ Clustering $-0.416$ Transitivity $-0.243$
PCA 4 $1.82\times 10^{-02}$ Density $-0.132$ Transitivity $-0.108$ Diameter $-0.088$
       PCA 0        PCA 1        PCA 2
Untrained
Refer to caption Refer to caption Refer to caption
Chem
Refer to caption Refer to caption Refer to caption
Social
Refer to caption Refer to caption Refer to caption
All
Refer to caption Refer to caption Refer to caption
Figure 4: PCA components of the encodings of each model across the whole validation dataset, scattered against the most correlated metrics, as in Table 4.
A.3 Loss Curves
Appendix B Further Results
B.1 Linear Transfer
Table 5: Scores for linear models applied to the encodings of each validation dataset, with and without node labels. For the classification datasets (ogbg-molclintox and Twitch Egos) we report AUROC, other datasets report RMSE. Some datasets, where results are very similar for each model, are excluded. Bold text indicates a superior result. For clarity the results for a given dataset are scaled by a constant factor.
Untrained Chem Social All Magnitude
molfreesolv 4.21 $\pm$ 4.13 3130 $\pm$ 5990 2.00 $\pm$ 0.869 1.01 $\pm$ 1.02 $10^{3}$
molesol 48.3 $\pm$ 90.2 817 $\pm$ 1150 48.1 $\pm$ 62.9 28.1 $\pm$ 38.1 $10^{0}$
mollipo 29.4 $\pm$ 17.9 14.9 $\pm$ 2.0 4.45 $\pm$ 0.62 5.33 $\pm$ 0.56 $10^{0}$
molclintox 0.487 $\pm$ 0.005 0.440 $\pm$ 0.007 0.500 $\pm$ 0.000 0.500 $\pm$ 0.000 $10^{0}$
Facebook 10.6 $\pm$ 1.7 10.7 $\pm$ 20 7.73 $\pm$ 1.87 4.44 $\pm$ 5 $10^{-3}$
Twitch Egos 0.638 $\pm$ 0.012 0.610 $\pm$ 0.072 0.679 $\pm$ 0.002 0.694 $\pm$ 0.002 $10^{0}$
Roads 263 $\pm$ 0.5 256 $\pm$ 0.0 298 $\pm$ 0.3 2.85 $\pm$ 5 $10^{-5}$
Trees 7.66 $\pm$ 0.16 2.85 $\pm$ 0.0 2.98 $\pm$ 0.03 2.56 $\pm$ 0.05 $10^{-3}$
Community 4.36 $\pm$ 0.05 4.25 $\pm$ 0.00 12.0 $\pm$ 2 4.46 $\pm$ 0.07 $10^{-5}$
Random 6.46 $\pm$ 0.7 3.27 $\pm$ 0.0 1.20 $\pm$ 0.2 3.38 $\pm$ 0.0 $10^{-4}$
Untrained-Labels Chem-Labels Social-Labels All-Labels
molfreesolv 480 $\pm$ 172 $(1.13\pm 0.64)\times 10^{6}$ 90900 $\pm$ 83300 11000 $\pm$ 5600 $10^{0}$
molesol 29.4 $\pm$ 7.8 18.5 $\pm$ 7.3 531 $\pm$ 292 34.1 $\pm$ 33.0 $10^{0}$
mollipo 7.33 $\pm$ 1.37 7.25 $\pm$ 1.01 4.02 $\pm$ 0.92 5.01 $\pm$ 0.53 $10^{0}$
molclintox 0.511 $\pm$ 0.022 0.497 $\pm$ 0.001 0.428 $\pm$ 0.001 0.526 $\pm$ 0.060 $10^{0}$
Facebook 14.6 $\pm$ 11.1 10.8 $\pm$ 1.1 7.82 $\pm$ 1.9 4.43 $\pm$ 6 $10^{-3}$
B.2 Transfer Learning
Table 6: Scores for fine-tuning models trained with random node dropping on each validation dataset, with (Model-labels) and without (Model) node labels. For the classification datasets (ogbg-molclintox and Twitch Egos) we report AUROC, other datasets report RMSE. Some datasets, where results are very similar for each model, are excluded. Bold text indicates a result beats the unsupervised (GIN) baseline. Underlined text indicates a superior result, down to three significant figures. GIN results are duplicated from Table 1.
GIN Chem Social All
molfreesolv 4.978 $\pm$ 0.479 3.838 $\pm$ 0.120 4.041 $\pm$ 0.116 4.157 $\pm$ 0.138
molesol 1.784 $\pm$ 0.094 1.897 $\pm$ 0.075 1.520 $\pm$ 0.066 1.357 $\pm$ 0.037
mollipo 1.143 $\pm$ 0.031 1.123 $\pm$ 0.105 1.015 $\pm$ 0.013 0.983 $\pm$ 0.015
molclintox 0.530 $\pm$ 0.042 0.469 $\pm$ 0.030 0.431 $\pm$ 0.002 0.530 $\pm$ 0.086
molbbbp 0.565 $\pm$ 0.043 0.497 $\pm$ 0.027 0.548 $\pm$ 0.037 0.567 $\pm$ 0.032
molbace 0.677 $\pm$ 0.053 0.740 $\pm$ 0.026 0.686 $\pm$ 0.055 0.679 $\pm$ 0.067
molhiv 0.355 $\pm$ 0.059 0.283 $\pm$ 0.001 0.635 $\pm$ 0.046 0.644 $\pm$ 0.039
Facebook 0.311 $\pm$ 0.065 0.261 $\pm$ 0.101 0.159 $\pm$ 0.015 0.145 $\pm$ 0.012
Twitch Egos 0.72 $\pm$ 0.006 0.460 $\pm$ 0.041 0.744 $\pm$ 0.006 0.743 $\pm$ 0.003
Roads 0.46 $\pm$ 0.073 0.228 $\pm$ 0.118 0.150 $\pm$ 0.015 0.132 $\pm$ 0.018
Trees 0.203 $\pm$ 0.012 0.415 $\pm$ 0.250 0.135 $\pm$ 0.015 0.125 $\pm$ 0.009
Community 0.743 $\pm$ 0.106 0.084 $\pm$ 0.083 0.080 $\pm$ 0.018 0.074 $\pm$ 0.021
Random 0.529 $\pm$ 0.228 0.123 $\pm$ 0.075 0.076 $\pm$ 0.010 0.058 $\pm$ 0.005
GIN-Labels Chem-Labels Social-Labels All-Labels
molfreesolv 4.854 $\pm$ 0.513 3.819 $\pm$ 0.136 4.113 $\pm$ 0.070 3.984 $\pm$ 0.162
molesol 1.707 $\pm$ 0.073 1.809 $\pm$ 0.139 1.480 $\pm$ 0.067 1.378 $\pm$ 0.031
mollipo 1.143 $\pm$ 0.031 1.182 $\pm$ 0.128 0.986 $\pm$ 0.017 0.997 $\pm$ 0.011
molclintox 0.497 $\pm$ 0.038 0.474 $\pm$ 0.036 0.430 $\pm$ 0.002 0.474 $\pm$ 0.056
molbbbp 0.561 $\pm$ 0.034 0.491 $\pm$ 0.032 0.543 $\pm$ 0.016 0.565 $\pm$ 0.036
molbace 0.698 $\pm$ 0.050 0.751 $\pm$ 0.024 0.677 $\pm$ 0.030 0.675 $\pm$ 0.088
molhiv 0.374 $\pm$ 0.052 0.286 $\pm$ 0.008 0.615 $\pm$ 0.047 0.628 $\pm$ 0.044
Facebook 0.749 $\pm$ 0.250 0.354 $\pm$ 0.166 0.151 $\pm$ 0.016 0.156 $\pm$ 0.022
Superior Result 19.0% 19.0% 19.0% 61.9%
Beats Baseline 47.6% 76.2% 90.5%
Table 7: Scores for fine-tuning models trained with random edge dropping on each validation dataset, with (Model-labels) and without (Model) node labels. For the classification datasets (ogbg-molclintox and Twitch Egos) we report AUROC, other datasets report RMSE. Some datasets, where results are very similar for each model, are excluded. Bold text indicates a result beats the unsupervised (GIN) baseline. Underlined text indicates a superior result, down to three significant figures. GIN results are duplicated from Table 1.
GIN Chem Social All
molfreesolv 4.978 $\pm$ 0.479 134.800$\pm$ 225.200 4.102 $\pm$ 0.049 4.197 $\pm$ 0.095
molesol 1.784 $\pm$ 0.094 14.050 $\pm$ 23.760 1.312 $\pm$ 0.094 1.571 $\pm$ 0.059
mollipo 1.129 $\pm$ 0.040 1.034 $\pm$ 0.029 0.989 $\pm$ 0.011 1.068 $\pm$ 0.026
molclintox 0.530 $\pm$ 0.042 0.449 $\pm$ 0.013 0.426 $\pm$ 0.006 0.461 $\pm$ 0.029
molbbbp 0.565 $\pm$ 0.043 0.570 $\pm$ 0.040 0.602 $\pm$ 0.025 0.603 $\pm$ 0.029
molbace 0.677 $\pm$ 0.053 0.768 $\pm$ 0.006 0.783 $\pm$ 0.045 0.756 $\pm$ 0.030
molhiv 0.355 $\pm$ 0.059 0.576 $\pm$ 0.152 0.647 $\pm$ 0.066 0.501 $\pm$ 0.135
Facebook 0.311 $\pm$ 0.065 3.70 $\pm$ 6.95 0.180 $\pm$ 0.019 0.237 $\pm$ 0.063
Twitch Egos 0.72 $\pm$ 0.006 0.717 $\pm$ 0.005 0.753 $\pm$ 0.006 0.739 $\pm$ 0.007
Roads 0.46 $\pm$ 0.073 0.156 $\pm$ 0.007 0.167 $\pm$ 0.039 0.252 $\pm$ 0.029
Trees 0.203 $\pm$ 0.012 0.143 $\pm$ 0.009 0.161 $\pm$ 0.019 0.187 $\pm$ 0.020
Community 0.743 $\pm$ 0.106 0.014 $\pm$ 0.000 0.146 $\pm$ 0.024 0.314 $\pm$ 0.069
Random 0.529 $\pm$ 0.228 0.055 $\pm$ 0.012 0.107 $\pm$ 0.019 0.289 $\pm$ 0.098
GIN-Labels Chem-Labels Social-Labels All-Labels
molfreesolv 4.854 $\pm$ 0.513 103.400$\pm$232.200 4.106 $\pm$ 0.021 4.163 $\pm$ 0.069
molesol 1.707 $\pm$ 0.073 8.006 $\pm$ 7.850 1.358 $\pm$ 0.028 1.597 $\pm$ 0.069
mollipo 1.143 $\pm$ 0.031 1.043 $\pm$ 0.026 0.995 $\pm$ 0.014 1.063 $\pm$ 0.038
molclintox 0.497 $\pm$ 0.038 0.446 $\pm$ 0.020 0.422 $\pm$ 0.009 0.475 $\pm$ 0.031
molbbbp 0.561 $\pm$ 0.034 0.535 $\pm$ 0.079 0.612 $\pm$ 0.015 0.589 $\pm$ 0.031
molbace 0.698 $\pm$ 0.050 0.763 $\pm$ 0.018 0.776 $\pm$ 0.032 0.744 $\pm$ 0.034
molhiv 0.374 $\pm$ 0.052 0.632 $\pm$ 0.087 0.652 $\pm$ 0.100 0.479 $\pm$ 0.102
Facebook 0.749 $\pm$ 0.250 2.80 $\pm$ 4.14 0.179 $\pm$ 0.020 0.263 $\pm$ 0.067
Superior Result 9.5% 19.0% 71.4% 4.8%
Beats Baseline 52.4% 90.5% 90.5%
Table 8: Scores for fine-tuning models pre-trained on non-molecules, each with one non-molecular pre-training dataset excluded, with (Model-features) and without (Model) features. For the classification datasets we report AUROC and for regression datasets RMSE. Some datasets, where results are very similar for each model, are excluded. Bold text indicates a result performs equally (or better) than the model pre-trained on molecules (Chem).
Social (by excluded pre-training dataset)
Chem Twitch Neurons Cora Facebook Roads
molfreesolv 3.89 $\pm$ 0.21 4.14 $\pm$ 0.07 4.09 $\pm$ 0.02 4.17 $\pm$ 0.03 3.97 $\pm$ 0.11 4.24 $\pm$ 0.16
molesol 1.44 $\pm$ 0.09 1.37 $\pm$ 0.08 1.64 $\pm$ 0.06 1.67 $\pm$ 0.04 1.58 $\pm$ 0.13 1.59 $\pm$ 0.07
mollipo 1.04 $\pm$ 0.02 1.00 $\pm$ 0.01 0.99 $\pm$ 0.01 1.01 $\pm$ 0.01 1.00 $\pm$ 0.01 1.01 $\pm$ 0.01
molclintox 0.51 $\pm$ 0.03 0.48 $\pm$ 0.12 0.44 $\pm$ 0.03 0.43 $\pm$ 0.00 0.43 $\pm$ 0.01 0.54 $\pm$ 0.10
molbbbp 0.54 $\pm$ 0.04 0.56 $\pm$ 0.04 0.40 $\pm$ 0.02 0.60 $\pm$ 0.02 0.62 $\pm$ 0.01 0.51 $\pm$ 0.02
molbace 0.63 $\pm$ 0.13 0.73 $\pm$ 0.04 0.62 $\pm$ 0.03 0.71 $\pm$ 0.05 0.70 $\pm$ 0.04 0.69 $\pm$ 0.02
molhiv 0.54 $\pm$ 0.13 0.42 $\pm$ 0.06 0.65 $\pm$ 0.07 0.54 $\pm$ 0.15 0.66 $\pm$ 0.04 0.61 $\pm$ 0.06
Facebook 1.17 $\pm$ 0.17 0.29 $\pm$ 0.02 0.14 $\pm$ 0.00 0.84 $\pm$ 0.01 0.14 $\pm$ 0.01 0.22 $\pm$ 0.06
Twitch Egos 0.69 $\pm$ 0.02 0.73 $\pm$ 0.03 0.74 $\pm$ 0.01 0.73 $\pm$ 0.01 0.74 $\pm$ 0.01 0.77 $\pm$ 0.01
Roads 0.57 $\pm$ 0.07 0.19 $\pm$ 0.02 0.11 $\pm$ 0.00 0.12 $\pm$ 0.01 0.11 $\pm$ 0.00 0.18 $\pm$ 0.01
Trees 0.24 $\pm$ 0.03 0.15 $\pm$ 0.01 0.12 $\pm$ 0.00 0.13 $\pm$ 0.01 0.12 $\pm$ 0.00 0.15 $\pm$ 0.01
Community 0.72 $\pm$ 0.13 0.28 $\pm$ 0.03 0.02 $\pm$ 0.00 0.07 $\pm$ 0.01 0.02 $\pm$ 0.01 0.22 $\pm$ 0.03
Random 0.59 $\pm$ 0.09 0.15 $\pm$ 0.01 0.05 $\pm$ 0.00 0.09 $\pm$ 0.02 0.04 $\pm$ 0.00 0.13 $\pm$ 0.02
Social-Features (by excluded pre-training dataset)
Chem Twitch Neurons Cora Facebook Roads
molfreesolv 4.16 $\pm$ 0.29 3.30 $\pm$ 0.32 3.90 $\pm$ 0.24 3.82 $\pm$ 0.31 3.67 $\pm$ 0.37 3.83 $\pm$ 0.38
molesol 1.62 $\pm$ 0.31 1.03 $\pm$ 0.08 1.03 $\pm$ 0.07 1.26 $\pm$ 0.16 1.13 $\pm$ 0.16 1.17 $\pm$ 0.10
mollipo 1.09 $\pm$ 0.06 0.93 $\pm$ 0.03 0.83 $\pm$ 0.03 0.92 $\pm$ 0.05 0.81 $\pm$ 0.03 0.94 $\pm$ 0.02
molclintox 0.52 $\pm$ 0.03 0.49 $\pm$ 0.03 0.76 $\pm$ 0.13 0.72 $\pm$ 0.10 0.76 $\pm$ 0.18 0.74 $\pm$ 0.06
molbbbp 0.61 $\pm$ 0.03 0.64 $\pm$ 0.02 0.64 $\pm$ 0.02 0.61 $\pm$ 0.03 0.62 $\pm$ 0.02 0.64 $\pm$ 0.03
molbace 0.70 $\pm$ 0.07 0.70 $\pm$ 0.04 0.69 $\pm$ 0.07 0.68 $\pm$ 0.05 0.67 $\pm$ 0.06 0.72 $\pm$ 0.05
molhiv 0.37 $\pm$ 0.07 0.41 $\pm$ 0.06 0.55 $\pm$ 0.11 0.48 $\pm$ 0.13 0.63 $\pm$ 0.07 0.52 $\pm$ 0.11
Table 9: Results for fine-tuning for three different layer architectures (Graph Convolution Network, Graph Attention Network, Graph Isomorphism Network) and pre-training data split on molecular benchmark datasets with node and edge features included. On regression datasets (freesolv, esol, lipo) we report MSE, and on classification datasets (bbbp, clintox, bace, hiv) we report 1 - AUROC.
GCN GAT GIN
None Chem Social All None Chem Social All None Chem Social All
Best 2.46 2.38 1.87 2.29 2.55 3.87 3.45 3.34 3.22 3.66 2.98 2.00
freesolv Mean 2.92 3.01 2.62 2.86 3.02 4.02 3.92 4.16 4 4.16 3.61 3.01
$\pm$ Dev. 0.309 0.456 0.363 0.38 0.292 0.113 0.271 0.549 0.434 0.288 0.428 0.594
Best 1.32 0.881 0.896 1.00 1.15 1.40 1.07 1.08 1.48 1.25 0.856 0.807
esol Mean 1.507 0.977 0.965 1.10 1.33 1.69 1.49 1.93 1.72 1.62 1.03 0.897
$\pm$ Dev. 0.095 0.072 0.055 0.094 0.116 0.192 0.262 0.636 0.245 0.306 0.165 0.042
Best 1.02 0.745 0.739 0.779 0.897 0.911 0.979 0.936 1.02 1.05 0.759 0.759
lipo Mean 1.14 0.782 0.765 0.827 1.11 1.03 1.06 1.03 1.09 1.09 0.798 0.788
$\pm$ Dev. 0.066 0.016 0.027 0.042 0.097 0.05 0.093 0.064 0.053 0.056 0.023 0.016
Best 0.303 0.329 0.318 0.327 0.324 0.329 0.322 0.369 0.357 0.347 0.334 0.290
bbbp Mean 0.335 0.375 0.352 0.387 0.352 0.433 0.391 0.416 0.399 0.391 0.374 0.335
$\pm$ Dev. 0.025 0.029 0.02 0.029 0.022 0.058 0.044 0.041 0.03 0.027 0.026 0.021
Best 0.0783 0.150 0.0879 0.0663 0.0471 0.5 0.5 0.489 0.112 0.444 0.106 0.0543
clintox Mean 0.126 0.532 0.305 0.468 0.157 0.5 0.5 0.498 0.351 0.485 0.138 0.164
$\pm$ Dev. 0.036 0.135 0.201 0.134 0.058 0.002 0 0.003 0.124 0.025 0.023 0.054
Best 0.177 0.239 0.195 0.200 0.156 0.238 0.268 0.270 0.210 0.243 0.232 0.203
bace Mean 0.233 0.292 0.28 0.27 0.243 0.378 0.371 0.339 0.247 0.305 0.304 0.276
$\pm$ Dev. 0.036 0.038 0.067 0.048 0.039 0.085 0.08 0.061 0.025 0.073 0.05 0.048
Best 0.466 0.255 0.249 0.245 0.536 0.286 0.308 0.283 0.611 0.533 0.278 0.272
hiv Mean 0.589 0.371 0.345 0.352 0.63 0.514 0.397 0.509 0.679 0.626 0.383 0.434
$\pm$ Dev. 0.074 0.064 0.119 0.072 0.063 0.166 0.082 0.15 0.032 0.066 0.106 0.131
B.3 Node and Edge Transfer
As in graph-level transfer we replace the original input head with a three-layer MLP for both node classification and edge prediction. For node classification, we use a single-layer MLP as an output head. For edge prediction, the embeddings from the encoder are concatenated for the relevant nodes, then passed to a single-layer MLP.
For edge prediction we sample the two-hop ego network of each node from the original graph. For node prediction we take the three-hop ego network, but limit the scale to five neighbours per node. As larger graphs we employ the Facebook and Cora graphs, as well as the Citeseer and PubMed graphs from the same paper. For edge prediction we also include the Amazon Computers and Photo graphs from Shchur et al. [30]. For node classification we also employ the GitHub and LastFM Asia graphs from Rozemberczki et al. [25]. For edge prediction we take 5% splits for training and testing edges. For node classification we take a 20% split. The results of this transfer can be found in Table 10.
Table 10: Scores for fine-tuning the complete model with node and edge features for node classification and edge prediction. For both node classification and edge prediction we report accuracy. Formatting is the same as Table 1.
Edge Pred. GIN-Features All (CL)-Features Chem-Features Social-Features All-Features
Amazon Comp. 0.85 $\pm$ 0.06 0.82 $\pm$ 0.08 0.84 $\pm$ 0.04 0.86 $\pm$ 0.05 0.86 $\pm$ 0.03
Amazon Photo. 0.86 $\pm$ 0.04 0.86 $\pm$ 0.05 0.89 $\pm$ 0.03 0.88 $\pm$ 0.05 0.91 $\pm$ 0.03
Citeseer 0.81 $\pm$ 0.01 0.78 $\pm$ 0.02 0.75 $\pm$ 0.02 0.75 $\pm$ 0.01 0.80 $\pm$ 0.01
PubMed 0.97 $\pm$ 0.01 0.98 $\pm$ 0.01 0.96 $\pm$ 0.02 0.98 $\pm$ 0.01 0.97 $\pm$ 0.01
Node Class. GIN-Features - Chem-Features Social-Features All-Features
Citeseer 0.75 $\pm$ 0.01 0.71 $\pm$ 0.05 0.72 $\pm$ 0.04 0.72 $\pm$ 0.01 0.74 $\pm$ 0.01
GitHub 0.89 $\pm$ 0.00 0.89 $\pm$ 0.02 0.89 $\pm$ 0.00 0.88 $\pm$ 0.02 0.89 $\pm$ 0.00
LastFM 0.79 $\pm$ 0.00 0.76 $\pm$ 0.02 0.78 $\pm$ 0.01 0.79 $\pm$ 0.01 0.79 $\pm$ 0.00
PubMed 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00
Table 11: Our results for the ToP-All and ToP-Social models, as well as a fully supervised baseline, presented alongside the same results from other pre-training works. Results for other works are taken verbatim from the respective work. On regression datasets (freesolv, esol, lipo) AD-GCL uses a downstream linear model instead of fine-tuning. Additionally these other works pre-train across Zinc2M [8], a dataset of 2 million molecules with all features included, which represents significantly more data than our amalgamated dataset. Notably these works report results for the pre-training epoch that produced the best results on these downstream datasets.
GIN InfoGraph [36] GraphCL [13] AD-GCL [32] ToP-Social ToP-All
freesolv 4.00 $\pm$ 0.43 - - 5.15 $\pm$ 0.62 3.61 $\pm$ 0.43 3.01 $\pm$ 0.59
esol 1.72 $\pm$ 0.25 - - 1.22 $\pm$ 0.09 1.03 $\pm$ 0.17 0.90 $\pm$ 0.04
lipo 1.09 $\pm$ 0.05 - - 0.84 $\pm$ 0.03 0.80 $\pm$ 0.02 0.79 $\pm$ 0.02
clintox 0.65 $\pm$ 0.12 0.70 $\pm$ 0.03 0.76 $\pm$ 0.03 0.80 $\pm$ 0.04 0.86 $\pm$ 0.02 0.84 $\pm$ 0.05
bbbp 0.60 $\pm$ 0.03 0.69 $\pm$ 0.01 0.70 $\pm$ 0.01 0.70 $\pm$ 0.01 0.62 $\pm$ 0.03 0.67 $\pm$ 0.02
bace 0.75 $\pm$ 0.03 0.76 $\pm$ 0.02 0.75 $\pm$ 0.01 0.79 $\pm$ 0.01 0.70 $\pm$ 0.05 0.72 $\pm$ 0.05
hiv 0.32 $\pm$ 0.03 0.76 $\pm$ 0.02 0.78 $\pm$ 0.01 0.78 $\pm$ 0.01 0.62 $\pm$ 0.11 0.57 $\pm$ 0.13