### v4

Abstract
The principal benefit of unsupervised representation learning is that a pre-trained model can be fine-tuned where data or labels are scarce. Existing approaches for graph representation learning are domain specific, maintaining consistent node and edge features across the pre-training and target datasets. This has precluded transfer to multiple domains. We present Topology Only Pre-Training (ToP), a graph pre-training method based on node and edge feature exclusion. We show positive transfer on evaluation datasets from multiple domains, including domains not present in pre-training data, running directly contrary to assumptions made in contemporary works. On 75% of experiments, ToP models perform significantly ($P\leq 0.01$) better than a supervised baseline. Performance is significantly positive on 85.7% of tasks when node and edge features are used in fine-tuning. We further show that out-of-domain topologies can produce more useful pre-training than in-domain. Under ToP we show better transfer from non-molecule pre-training, compared to molecule pre-training, on 79% of molecular benchmarks. Against the limited set of other generalist graph models ToP performs strongly, including against models with many orders of magnitude larger. These findings show that ToP opens broad areas of research in both transfer learning on scarcely populated graph domains and in graph foundation models.
1 Introduction
Representation learning allows models to convert the useful information from rich data, for example text and images, into flexible vector encodings. As pre-training, then fine-tuning on downstream tasks, representation learning leads to improved performance [6, 22, 9]. This is an invaluable technique where downstream labels are scarce, allowing expressive models to be applied where otherwise performance would be unsatisfactory. Graph representation learning extends these methods to structured data, allowing much work on unsupervised learning over large datasets of graphs from a given domain [16, 5, 21].
In other domains, such as images and text, it is common to pre-train on a large amalgamated dataset of samples, aiming to maximise coverage of the space [38, 45, 33]. Large models pre-trained on these massive datasets then show strong performance on a large range of downstream tasks. Models with this combination of large size, large data and strong performance are commonly referred to as foundation models [30].
There are substantial obstacles to taking this approach for graph data, the largest being feature heterogeneity between domains. Images or text from multiple domains are fundamentally the same data-type, uniform-size arrays of pixel values or sequences of tokens respectively, meaning that the multi-domain problem can be tackled head-on through data and model scale. Graphs, on the other hand, have a large diversity of possible feature sets. We present model examples of feature sets for social networks and molecules in Figure 1. How a single model can handle highly heterogenous feature sets like these, for example both a text biography and the atomic number of an atom, is an open and challenging problem.
Figure 1: Graph data has rich, highly varied feature sets between domains. Here we provide model examples from social media (left) and chemistry (right). Social media feature sets might include a combination of textual, image and numeric types. Chemical feature sets are primarily numeric and categorical, representing properties of component atoms that might be difficult to infer for deep-learning models.
This has lead current research to argue or assume, but not show, that multi-domain pre-training on graphs is not feasible [26, 42, 56]. As a consequence there are few works attempting learning across multiple domains of graphs. Some very recent works have proposed approaches using LLMs as feature transformations [40, 55, 52, 25], effectively moving a source domain into a unified text-featured domain. This drastically increases computational expense, and when features are not easily or commonly expressible as text, for example continuous features, LLM approaches presumably cannot be used.
In downstream tasks feature information is often essential, so generalised approaches should allow them to be included in downstream applications, unlike in the very limited set of approaches for generalised graph pre-training [32]. The lack of a generalised pre-training method, which allows domain features to be included downstream, excludes areas of data sparsity or scarcity from the benefits of pre-training. Even on areas where data is plentiful enough for pre-training, it requires a different pre-trained model to be used each time, increasing the time, complexity, and cost of usage. Multi-domain pre-training would avoid these issues, both opening transfer learning as an option for domains with limited access to data, and providing a baseline model on which applications can be built.
1.1 Pre-Training without Features
Graph learning can rely on information from both structure and features, with some of that information coming only from structure. Therefore, beneficial learning during pre-training can occur with features excluded. Motivated by this view we introduce Topology Only Pre-Training (ToP), a generalised pre-training method for graph models. Under ToP pre-training is conducted with node and edge features excluded, but during downstream application they can be re-included by swapping components of a given model.
We investigate ToP as a pre-training method for graph data that is not limited to a given domain or task. The result of such pre-training is a single model that can be transferred with positive performance increases on arbitrary graph-level tasks and domains. Experimentally we introduce and test two hypotheses:
Hypothesis 1: Excluding Features
Pre-training on multiple domains with node and edge features excluded can lead to consistent positive transfer compared to fully-supervised, task-specific models.
Hypothesis 2: Sample Diversity
With the exclusion of node and edge features out-of-domain pre-training can have greater benefits than in-domain pre-training across downstream tasks.
Findings
We evaluate ToP using graph constrastive learning with both random and adversarial augmentations. This evaluation includes the construction of an amalgamated, multi-domain dataset, spanning from road networks to molecules. We pre-train multiple models on different subsets of this amalgamated dataset and with varied augmentation schemes. We perform multiple fine-tuning runs for each model in order to show statistical significance for our experiments. Through this evaluation we contribute the following empirical findings:
  1. 1.
    Effective pre-training. Pre-training on varied domains, under the exclusion of node and edge features, leads to positive transfer. We find that under fine-tuning, results are at worst on-par, and on 76% of tasks significantly better, than a non-pretrained baseline.
  2. 2.
    Increased performance from domain diversity. Pre-training without features from varied non-target domains, instead of the target domain, can bring greater downstream performance benefits. In the absence of features in pre-training a non-molecule learner consistently produces positive transfer compared to a molecule learner. On 56% of tasks the non-molecule model performs significantly better, and never significantly worse, than the molecule-only model.
Contributions
We present two main contributions, although the work contains several more.
Our first contribution is ToP, a multi-domain pre-training method for graph models, with original domain features included during downstream tasks. Therefore, and under recent definitions, models pre-trained using ToP are close to being foundation graph models [2, 25, 26, 56]. ToP is the first method that, to our knowledge, has the following desirable properties:
  1. 1.
    Positive transfer across multiple, highly diverse domains
  2. 2.
    Original node and edge features during downstream transfer
  3. 3.
    No very large auxiliary models
  4. 4.
    No downstream feature transformations
Across molecular benchmarks our implementation of ToP out-performs benchmarks from the limited set of generalist GNN approaches that exist. Despite not aiming to out-perform domain specific specialist models, ToP performs better-than or on-par with generalist LLMs many orders of magnitude larger in both parameters and training dataset size. This provides great utility when pre-training data is not abundant.
Our second contribution is a new research direction: with domain features excluded, diverse non-domain pre-training can be more valuable than in-domain. This runs contrary to the single-domain training paradigm, and directly opposite to the assumptions made in other works [28, 42]. In Mao et al [28]’s review of LLMs for for multi-domain graph models, enumerating why they show promise, it is assumed that “…training on [graphs from multiple domains] simultaneously shows no positive transfer benefit while increasing the risk of the negative transfer”. Our work shows that the opposite is true when pre-training is conducted only on structure. This implies that greater performance through ToP can be achieved simply by increasing the scale and diversity of the pre-training dataset.
Organisation
In Section 2, we describe the notation and concepts necessary to understand our method and provide an overview of related works. In Section 3, we describe our generalised pre-training method ToP, and our hypotheses about it as a method. In Section 4, we describe how we evaluate the hypotheses detailed above, and present the findings of those experiments. In Section 5, we discuss the implications of the results from these experiments, as well as list the assumptions we make and threats to validity. In Section 6, we conclude with avenues for future research.
2 Background and Related Work
Here we describe the concepts necessary to understand the rest of our method, as well as closely related works.
Representation Learning
In most contexts, representation learning is the process of forming useful numerical encodings of non-tabular input data through an ‘encoder’ [1]. Let $X\in\mathcal{X}$ represent a random variable $X$ in the data space $\mathcal{X}$, and $Y\in\mathcal{Y}$ the same for the output space. Representation learning aims, given their joint distribution $P(X,Y)$, to learn a mapping $X\rightarrow\hat{X}$ that maximises their mutual information $I(\hat{X};Y)$. In basic terms, representation learners aim to transfer some piece of rich data, for example images, into a vector that describes that piece of data.
Contrastive Learning (CL)
CL is a semi-supervised school of representation learning based on altering input samples [22]. Let $\mathcal{T}$ be a set of small transformations called ‘augmentations’, where $t\in\mathcal{T}:X\rightarrow X^{\prime}$. These augmentations are generally designed to manipulate the input sample in a way non-destructive to the ‘useful’ information it represents. Augmentations can be understood as a method to increase the semantic coverage in the input space. Common choices for images, for example, are gray-scaling and rotations. Intuitively, the augmentation defines an invariant transformation for the encoder to learn. Given an encoder’s mapping $F(X_{i})\rightarrow\hat{X_{i}}$, multiple augmentations of the same input sample should be as similar as possible. Again in simple terms, contrastive learners aim to represent slight variations on the same sample in similar ways, for example through cosine similarity.
Graph Representation Learning
Hamilton [15] provides a comprehensive summary of methods for graph representation learning. Traditional methods to encode graphs as vectors include graph statistics such as centrality measures and clustering coefficients, but also kernel methods such as the Weisfeiler-Lehman kernel [39]. Expressive Graph Neural Networks (GNNs) models, generally through parameterised message-passing, can produce powerful representations of graphs. A number of GNN architectures have been proposed with the Graph Isomorphism Network (GIN) being the most expressive in message-passing representation learning contexts [59].
Graph Contrastive Learning
Multiple studies exist for graph contrastive learning [8, 14, 17, 20, 24, 54, 60]. The typical graph contrastive learning formulation, Graph Contrastive Learning (GraphCL), is put forward by You et al [54]. In GraphCL and many subsequent works, such augmentations are performed randomly using fixed parameters, i.e., ‘drop 20% of nodes’. These augmentation strategies do not vary based on the graph-instance being augmented.
Adversarial Graph Contrastive Learning
The quality of augmentations is critical for graph contrastive learning [54, 60]. To maximise expressivity in these augmentations adversarial strategies have been proposed [17, 57]. Suresh et al [43] propose Adversarial-Graph Contrastive Learning (AD-GCL), with their method training both an encoder and a ‘view-learner’ $\mathcal{T}_{\theta}$. AD-GCL performs the standard GCL protocol with weight updates to a parameterised ($\theta$) set of augmenting functions $\mathcal{T_{\theta}}$ through the contrastive loss. Here $\mathcal{T_{\theta}}$ learns augmentations that minimise mutual information between the source and augmented graphs through edge dropping. The encoder, adversarially to this, learns to maximise shared information between the respective augmentations. The updates for $\mathcal{T_{\theta}}$ are regularised such that augmentations are not entirely destructive.
Large Language Models on Graphs
Large Language Models (LLMs) are models pre-trained, typically generatively, across massive corpuses of natural language data. Relying on scaling through both dataset size and model size, such models are highly performant on natural language processing. The first foundation models, though trained in the same manner, were smaller and typically required fine-tuning for strong performance on a specific task. Bi-Directional Encoder Representations from Transformers (BERT) is the archetypal example here [9]. The most recent LLMs vary in size from a few billion parameters [10] to many hundreds of billions [33].
LLMs have in very recent research been applied to building multi-domain graph models, often termed Graph Foundation Models (GFMs) in their respective works. Some have argued that LLMs themselves, by default, may be GFMs [28]. LLMs carry the significant benefit of allowing one-shot or few-shot application on downstream tasks in some circumstances. However, given their size, LLMs can drastically increase the cost and complexity of using graph models.
The simplest approach for applying LLMs on graph data is simply to encode a graph structure in natural language, along with features. We model our description of this approach on Fatemi et al [12]. Here a graph structure is encoded via some schema, for example the edge $(Callum,Zack)$ might be encoded ‘Callum and Zack are connected’ in a social network. Fatemi et al [12] find that the choice of encoding schema, for example ‘is friends with’ vs ‘is connected to’, has a significant impact on performance, and that in-general this approach is limited.
Alternatively the performant capacity of LLMs as encoders has been used to unify the graph feature space across domains. By describing the features of a node or edge in natural language, the LLMs encodings of that new textual feature can be shared between domain feature sets [40, 55, 52, 25]. This assumes and requires that:
  1. 1.
    Original features are easily encoded in text
  2. 2.
    Natural language versions of features are meaningful
  3. 3.
    LLM encodings are in fact a coherent embedding space
For natural-language features, the approach is trivial, and numeric features are converted to text via a schema. An example for molecular graphs would be ‘An atom with X protons, Y covalent bonds, which (is/is not) part of a ring structure’. These textual features are then passed to an LLM to form numeric encodings as features. This has the significant advantage of allowing one-shot or few-shot downstream use, assuming that downstream text features are close to in-domain. The approach has yet to be applied to graphs with rich feature sets. We assume this because it is not clear how previous approaches would encode non-intuitive graph features textually, such as continuous ones. Many of these initial works use multiple domains where features are already text, and do not attempt to move other feature sets into textual features.
Another approach has emerged in LLMs for chemistry, where there are expressive sequential descriptions of molecular graphs commonly used in natural language [27, 58, 4]. Here LLM approaches use sequential SMILES encodings of molecules, for example O=C=O from CO2, to represent molecules at some point of their pipeline. SMILES strings, as a natural language expression of a graph structure, are a useful tool for LLM approaches. Most other graph structures, for example social networks, are much harder to express in natural language. A core assumption of these models, with strong evidence for validity through experimentation, is that forming datasets of text surrounding these SMILES encodings using chemical literature allows learning about chemical dynamics to occur. Where these models do not include explicit usage of graph structures, for example InstructMol [4] and ChemDFM [58], they require far larger models and datasets for the same performance.
ToP carries few of the downsides associated with LLMs for graph data, yet achieves many of the same goals. We show that ToP models have strong downstream performance across domains on structure-only tasks, and in comparison, LLMs perform far more weakly. Despite the suitability of molecular tasks for LLMs, as described above, ToP performs at a level comparable to the best-performing LLMs on molecular tasks. This comes at a massively smaller time and computational cost compared to LLMs. Our reasonably large GNN encoders complete pre-training in less than half a day, on consumer hardware, with fine-tuning and inference both feasible and fast on consumer laptops.
Graph Constrastive Coding
A similar work to ToP is Graph Contrastive Coding (GCC) [32]. GCC begins by generating subgraphs from various large graphs through ego sampling. As augmentations it used random walks with restarts, with embeddings of the seed node used for its contrastive loss, resulting in a local-focus model. This is in contrast to this work, where we focus on individual graph level representations, instead of pre-training as a route to tasks on large graphs.
As in this work, GCC’s primary aim is to sidestep the issue of varied feature sets between graph domains. ToP simply removes feature sets during pre-training. GCC goes a step further, replacing existing node and edge features with positional information, principally the graph Laplacian. While this has obvious merit in encoding structural information, models can learn from specific local-level features, allowing overfitting and other feature-space biases to occur. Our assumption, validated through experiment, is that such an approach also makes re-introduction of domain-specific features much less effective. Where we do not extend GCC by including domain features during transfer, using structural encodings as in the original work, performance on out-of-domain data is weak. On out-of-domain structure-only tasks, ToP models perform more strongly or on-par with GCC models, despite lacking the advantage of structural features included with GCC.
3 ToP: Topology Only Pre-Training
Assume a dataset with features $X\in\mathbf{R}^{|V|\times D}$ and similar for edges, topology as an edgelist $E:\{(v_{1},v_{2}),\ldots\}$, and graph labels $y$, with graphs $G:\{X,E\}$. $y$ is some target label. When a given model aims to learn some mapping $f(G)\rightarrow y$, in reality this is $f(X,E)\rightarrow y$. The usual graph learning assumption is that topology $E$ and features $X$ must be considered together for information to be useful. In other words their combination has mutual information with a target $y$, $I(y;X,E)$.
We can assume that in the absence of features, structural information is still present, i.e. $I(y;E)\geq 0$. From this view, pre-training without features is well-motivated, with models learning generalised information about graph structures before features are introduced. This information from $E$ is useful for downstream tasks, even while missing task-specific information from features. This is an abstraction away from feature-specific correlations, with abstraction a core quality of effective representation learners [22]. Removing features $X$ necessarily removes their own information, but also the interdependence of features and structure:
$I(y;E,X)\geq I(y;X)+I(y;E)$ (1)
In situations where data is abundant, therefore, pre-training on both $X$ and $E$ will likely still to higher performance.
Graph without features are sampled from the same high dimensional space, which is highly constrained in real world graphs. As such they can be similar across highly varied domains, much as in the image space, even if they are entirely heterogenous with features included. On the local scale at which message passing GNNs operate, the same patterns may occur across any number of domains. In this case a wider range of topologies could span a wider range of the basic graph space, and so contain more useful information, than a narrow range closer to a downstream domain.
This pre-training on only topologies would give an initial bias towards structural features during downstream transfer. Such a structural bias is encouraged in other works, in particular those works on Topological GNNs [18, 7, 48]. Here layers or models themselves are designed to encourage learning from patterns in structure over patterns in features.
Topology Only Pre-Training (ToP) is a method for generalised pre-training of graph models. Under ToP pre-training is conducted only on topologies, with node and edge features excluded through replacement with identical values. This use of only topologies $G$, instead of graphs with specific features $\underline{G}$, allows pre-training on a multi-domain dataset. In the works prior to ToP, the use of such a graph dataset was not tractable [26, 42, 56].
Figure 2 shows a schematic for ToP models. ToP’s implementation is elegant: Features are removed from nodes and edges and replaced with some arbitrary identical value. In our case all nodes and edges carry the integer value 1. With the input head some simple MLP, the choice of value has no significant impact, as the GNN layers are presented the same vector for all nodes and edges. An encoder is then pre-trained using some unsupervised representation learning method. In our implementation we use contrastive learning for pre-training, given its strong performance in representation learning on graphs.
By swapping the input head of the pre-trained model, arbitrary node and edge features can be included during transfer, allowing integration of feature information alongside structural during downstream transfer. In the same manner the output head of a model can be replaced or excluded entirely, allowing transfer onto node and edge-level tasks.
Figure 2: A schematic for ToP models. During pre-training, features are replaced with a single integer label for each node or edge (all identical), which are then passed to a single-layer MLP. This results in an identical input vector for each node and edge. During transfer, a different input head is used, moving the original dimensions to the hidden dimensionality of the GNN encoder block. This allows arbitrary node and edge features to be included in transfer. Omitting the output head of the model allows transfer onto node and edge-level tasks.
3.1 Hypothesis 1: Excluding Features
Pre-training on multiple domains with node and edge features excluded can lead to consistent and positive transfer compared to fully-supervised, task-specific models. The motivating concept behind ToP is that as graph learning requires learning on both structural and feature levels, pre-training on only structure can lead to useful knowledge transfer. As such, a model that hasn’t undergone such pre-training should perform worse on downstream tasks when applied under the same supervised training conditions.
Supporting Hypothesis 1 means producing a model that has learnt useful representations of graphs in general. Our first step to achieving this is to construct a dataset that covers a variety of graph domains. Such a dataset should contain graphs that present a large diversity of multi-node patterns. Our approach is simple: include graphs from several domains and assume that there is adequate coverage of the multi-node patterns that occur in real-world graphs for a given set of downstream tasks. The same approach has been taken in the domains of text [3] and images [34], where enormous numbers of data samples are scraped from the web for generative pre-training. This relies on two assumptions 1) this dataset adequately covers the relevant space for downstream transfer, and 2) a single model can usefully represent such a space. These assumptions, as in other fields, are justified through performance metrics on downstream tasks on seen and unseen domains.
We formulate our test for Hypothesis 1, that pre-training on graphs without node or edge features is beneficial, as beating a simple supervised baseline. Said baseline is a model with the same architecture as the pre-trained model but without pre-training. Both models are then trained, or fine-tuned for the pre-trained model, on the same downstream datasets under the same conditions. On each downstream task, should the pre-trained model match or out-perform the supervised model, this represents positive transfer. Positive transfer on a task indicates that at least some of the learning during pre-training was beneficial to subsequent learning during fine-tuning. A positive transfer rate across datasets of more than 50% indicates that in-general transfer is positive, thus meeting the minimum criteria for Hypothesis 1. A high positive transfer rate strongly supports Hypothesis 1.
3.2 Hypothesis 2: Sample Diversity
With the exclusion of node and edge features, out-of-domain pre-training can have greater performance benefits than in-domain pre-training across downstream tasks. If structure-only learning is useful, as proposed by Hypothesis 1, it follows that out-of-domain of pre-training data with greater diversity may be more useful than in-domain. This relies on the assumption that a given encoder is learning to represent local node patterns, and as diverse data presents a wider range of such patterns, the model must learn more expressively how to represent such patterns.
From Hypothesis 2 we expect that for a given task models trained on a more diverse pre-training domain can bring better transfer than a model trained on the target domain. We do not argue that this is always true, only that in some settings out-of-domain samples can be more useful for pre-training than in-domain. This has clear parallels to the performance-by-scale approach taken in other fields [38, 9, 30]. To this end we compare two models, the first trained on a dataset from a single domain, as is the current practice. This single domain matches that of the downstream target dataset(s). The second model is instead trained on the rest of the data while excluding the target domain.
The total size of the pre-training datasets for each model should match for fair comparison. By pre-training both models under the same controlled conditions (hyperparameters etc.), then evaluating their performance under fine-tuning, we directly compare pre-training on a single domain to pre-training on multiple domains. General positive transfer from the multi-domain model, in comparison to transfer from the in-domain model, would constitute a clear supporting evidence for Hypothesis 2.
3.3 Models for Evaluation
In order to address these hypotheses, we apply ToP over five principal different combinations of pre-training data and augmentation scheme.
GIN
An untrained GIN model of identical architecture to the other methods. This serves to benchmark what constitutes positive transfer.
ToP-Chem
An AD-GCL model fit only on molecules. This constitutes an in-domain model, albeit in the absence of domain features.
ToP-Social
An AD-GCL model fit only on non-molecules. This serves as a contrast to ToP-Chem, allowing us to assess the benefits of varied out-of-domain pre-training in comparison to in-domain pre-training.
ToP-All
An AD-GCL model fit on all the training data. This is our primary measure of the effectiveness of ToP.
ToP-All (CL)
A model trained across the whole amalgamated dataset with the loss function from You et al [54] using random edge dropping ($\lambda=0.2$). This allows us to assess the impact of adversarial augmentations in ToP pre-training.
Models during pre-training, unless specified, are six GIN [53] layers with single linear layers as input and output. The GIN model acts as our supervised baseline (Hypothesis 1 in Section 4). The ToP-Chem/ToP-Social models indicate the impact of training data diversity on downstream transfer performance, particularly on our downstream molecular datasets (Hypothesis 2). The ToP-All (CL) model shows the impact of adversarial augmentations.
Figure 3: UMAP embeddings of encodings from each model, as well as an untrained GIN model. Triangular markers show molecular graphs, and circles non-molecules. We plot the centroid of each dataset for added clarity. Qualitatively the untrained and ToP-Chem models are noticeably more fragmented than the ToP-Social and ToP-All models. In turn the ToP-Social model is more fragmented than the ToP-All model.
4 Experiments
4.1 Datasets for Evaluation
Evaluating ToP requires collecting datasets of graphs from multiple domains. A portion of these datasets should have downstream tasks, the useful information for which relies on both structure and features, with varied weighting to one or the other. To this end we curate the following datasets, with Table 6 in the Appendix providing dataset statistics:
Pre-Training Only
molpcba [20, 50]
A curated dataset of 437,929 small molecules, from which we use 250,000. Used only in pre-training.
Cora [29]
A much used citation graph. Used only in pre-training.
Fly Brain [49]
The full neural connectome of a fruit fly larvae. As the original multi-graph is dense at 548,000 edges, we include a single edge between neurons if there are more than two synapses between them. Used only in pre-training.
Pre-Training and Evaluation
Facebook [37]
A single graph of page-page connections on Facebook. Used both in pre-training and downstream transfer. On a graph level the downstream task is predicting average clustering, and we also evaluate transfer to node classification. This task does not use feature information.
Twitch [35]
Ego networks from the streaming platform Twitch, used both in pre-training and downstream transfer, with a downstream binary classification task. This task does not use feature information.
Roads [23]
The road network of Pennsylvania. Used both in pre-training and downstream transfer, with the task set to be predicting graph diameter normalised by graph size. This task does not use feature information.
Evaluation Only
Benchmark Molecules [20]
molesol, molfreesolv, mollipo, molclintox, molbbbp, molbace, molhiv. Datasets sampled from MoleculeNet, with single-target classification or regression downstream tasks. These are used only in downstream transfer. These benchmarks have pre-prepared node and edge features, allowing us to measure the extent to which ToP models can integrate feature information alongside structural. We source train/validation/test splits from the Open Graph Benchmark [19].
Trees
Trees, of a randomly sampled maximum depth, and a randomly fixed probability of branching at each level, with a downstream task of predicting depth. Used only in downstream evaluation. This task does not use feature information.
Random
Erdos-Renyi (ER) graphs [11], with a number of nodes and edge probability randomly sampled, and predicting edge probabilities as a downstream task. Used only in downstream evaluation. This task does not use feature information.
Community
Small ER subgraphs, with a set intra-subgraph and random inter-subgraph connection probability, with inter-subgraph connectivity as target. Used only in downstream evaluation. This task does not use feature information.
Datasets from Single Graphs
As the Facebook, Cora, Roads and Fly Brain datasets consist of a single graph, we employ exploration sampling to construct datasets of many smaller graphs. We use a randomly selected exploration sampler for each sample [36], which in turn samples a random number of nodes in the range $24\leq|V|\leq 96$ from the source graph. The use of a selection of exploration samplers should further ensure that a large variety of graphs and structural patterns are included in the datasets.
Hyperparameters
We take the downstream task on each validation set where available and add their relevant scores (MSE and $1-$ AUROC) with a linear model as a crude monitor for the models’ representational abilities during pre-training. 5,000 samples are used for non-molecular datasets. We conduct a limited Bayesian hyperparameter sweep. The heuristic used is addition of evaluation scores on each dataset weighted by the number of samples in that dataset. Following the results of the hyperparameter sweep, we train each model for 100 epochs with a batch size of 512, learning rates of 0,001, embedding and projection dimensions of 300 and a regularisation constant of 0,2. This is very close to the default AD-GCL parameters. Figure 3 shows UMAP embeddings of encodings of the validation datasets for our models. Larger versions for the four ToP models can be found in the Appendix.
Compute
Pre-training was conducted on a workstation with an NVIDIA A6000, using around 16GB of VRAM capacity. Pre-training the ToP-All model took a little more than eight hours. Transfer results were produced on consumer laptops, primarily with an Apple M1 processor, and on GPU use around 3GB of video memory with batch sizes of 128.
4.2 Transfer
Here we apply our pre-trained models to various downstream tasks through transfer learning. This follows stages of abstraction away from ToP pre-training: first we fine tune on datasets without features, then on datasets with features included, and then compare ToP models to reasonable benchmarks. Finally we report performance on node and edge-level tasks.
Table 1: Scores for fine-tuning on each validation dataset. Molecular benchmarks here do not include node or edge features. For the classification datasets we report AUROC and for regression datasets RMSE. Bold text indicates a result beats the supervised (GIN) baseline. Underlined text indicates a superior result, down to three significant figures. Checkmarks ✓indicate that the ToP-All model significantly out-performs the GIN baseline at $p<0.01$. ‘Best’ indicates the proportion of datasets over which a model has superior performance, and ‘Positive’ the proportion of datasets over which performance is superior to the fully-supervised baseline.
GIN ToP-All (CL) ToP-Chem ToP-Social ToP-All
Facebook 0.31 $\pm$ 0.07 0.24 $\pm$ 0.06 1.17 $\pm$ 0.17 0.13 $\pm$ 0.00 0.13 $\pm$ 0.00
Twitch Egos 0.72 $\pm$ 0.01 0.74 $\pm$ 0.01 0.69 $\pm$ 0.02 0.74 $\pm$ 0.01 0.79 $\pm$ 0.01
Roads 0.46 $\pm$ 0.07 0.25 $\pm$ 0.03 0.57 $\pm$ 0.07 0.10 $\pm$ 0.00 0.10 $\pm$ 0.00
Trees 0.20 $\pm$ 0.01 0.19 $\pm$ 0.02 0.24 $\pm$ 0.03 0.12 $\pm$ 0.00 0.12 $\pm$ 0.00
Community 0.74 $\pm$ 0.11 0.31 $\pm$ 0.07 0.72 $\pm$ 0.13 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00
Random 0.53 $\pm$ 0.23 0.29 $\pm$ 0.10 0.59 $\pm$ 0.09 0.04 $\pm$ 0.00 0.04 $\pm$ 0.00
molfreesolv 4.98 $\pm$ 0.48 4.20 $\pm$ 0.10 3.89 $\pm$ 0.21 3.83 $\pm$ 0.11 4.26 $\pm$ 0.03
molesol 1.78 $\pm$ 0.09 1.57 $\pm$ 0.06 1.44 $\pm$ 0.09 1.47 $\pm$ 0.04 1.31 $\pm$ 0.03
mollipo 1.13 $\pm$ 0.04 1.07$\pm$ 0.03 1.04 $\pm$ 0.02 0.99 $\pm$ 0.01 0.97 $\pm$ 0.01
molclintox 0.53 $\pm$ 0.04 0.46 $\pm$ 0.03 0.51 $\pm$ 0.03 0.50 $\pm$ 0.05 0.55 $\pm$ 0.11
molbbbp 0.57 $\pm$ 0.04 0.60 $\pm$ 0.03 0.54 $\pm$ 0.04 0.60 $\pm$ 0.03 0.59 $\pm$ 0.02
molbace 0.68 $\pm$ 0.05 0.76 $\pm$ 0.03 0.63 $\pm$ 0.13 0.75 $\pm$ 0.03 0.71 $\pm$ 0.02
molhiv 0.36 $\pm$ 0.06 0.50 $\pm$ 0.14 0.54 $\pm$ 0.13 0.64 $\pm$ 0.05 0.66 $\pm$ 0.04
Best 0% 15.4% 0.0% 38.5% 69.2%
Positive 92.3% 38.5% 92.3% 100%
4.2.1 Features Excluded
We replace the output head with a simple two-layer MLP. We then perform 10 fine-tuning runs of 50 epochs on each validation dataset, with and without features, for each pre-trained model. Table 1 presents results of transfer and fine-tuning without features.
The ToP-All model consistently out-performs the other models under fine-tuning on the majority of datasets. At $p\leq 0.01$ the ToP-All model significantly out-performs the GIN baseline on 77% of datasets. When performance is not significantly better, it is never significantly worse. The standard deviation of the results on each dataset are lower than those for the non-pretrained model and the domain-specific pre-trained models. In Figure LABEL:fig:training-examples in the Appendix we show validation loss on the several datasets during fine-tuning for the ToP-All model and an un-pre-trained GIN with the same architecture. Here the advantages of pre-training are clear: the pre-trained model has consistently lower validation loss, seems to plateau later and at a lower loss, and has a lower deviation in that loss than the baseline GIN.
On our evaluation-only datasets (Trees, Community, Random), on which no models are pre-trained, we see markedly better performance from models that are pre-trained on a range of domains. In particular the ToP-All and ToP-Social models drastically out-perform our supervised baseline and ToP-Chem. On the Trees dataset, which intuitively and on aggregate measures is the closest of the evaluation-only datasets to molecular data, ToP-Chem also has negative transfer.
Pre-training only on molecules (‘Chem’ data subsets) results at-best on-par transfer compared to the supervised GIN baseline, with negative transfer on the majority of datasets. Pre-training only on non-molecules, in comparison, results in positive transfer on the large majority of datasets and pre-training methods. At $p\leq 0.01$ the ToP-Social model significantly out-performs the ToP-Chem model on the majority of datasets. As before, where performance is not significantly better, it is never significantly worse. This study corroborates the results presented in Table 1, with the Chem model consistently out-performed by non-molecular models. We find that under pre-training with random edge dropping performance actually drops when including molecular data (‘Social’ vs ‘All’). This suggests that AD-GCL is able to differentiate between molecules and non-molecules and applies different augmentations to these domains. Nonetheless, we show that using a variety of non-molecular graphs in-general provides a more solid pre-training foundation than including molecules for both methods of augmentation.
4.2.2 Features Included
We replace the input head of our model with a three layer Multi-Layer Perceptron (MLP) for both node and edge features, then perform fine-tuning runs on our molecular datasets. Table 2 presents the results of this transfer and fine-tuning. These tasks are assumed to require node and edge features to obtain good performance, and here positive transfer rates are similar to when no features are included. This shows that the focus on structure during ToP pre-training does not inhibit the use of feature-based information in downstream transfer. The ToP-All and ToP-Social models again demonstrate consistent positive transfer, and on most datasets their performance improves compared to when features were excluded. The same results are presented for different GNN backbones in Table 11 in the Appendix, but we perform no optimisation for these models, using hyperparameters from the GIN models here for the GAT and GCN models presented. As such these results are included only for completeness.
Table 2: Scores for fine-tuning pre-trained GIN models (and a non-pre-trained baseline) on the molecular evaluation datasets with node and edge features included via projection heads. Results are reported in the same manner as in Table 1.
GIN ToP-All (CL) ToP-Chem ToP-Social ToP-All
molfreesolv 4.00 $\pm$ 0.43 4.00 $\pm$ 0.24 4.16 $\pm$ 0.29 3.61 $\pm$ 0.43 3.01 $\pm$ 0.59
molesol 1.72 $\pm$ 0.25 1.18 $\pm$ 0.09 1.62 $\pm$ 0.31 1.03 $\pm$ 0.17 0.90 $\pm$ 0.04
mollipo 1.09 $\pm$ 0.05 0.99 $\pm$ 0.05 1.09 $\pm$ 0.06 0.80 $\pm$ 0.02 0.79 $\pm$ 0.02
molclintox 0.65 $\pm$ 0.12 0.56 $\pm$ 0.16 0.52 $\pm$ 0.03 0.86 $\pm$ 0.02 0.84 $\pm$ 0.05
molbbbp 0.60 $\pm$ 0.03 0.63 $\pm$ 0.04 0.61 $\pm$ 0.03 0.62 $\pm$ 0.03 0.67 $\pm$ 0.02
molbace 0.75 $\pm$ 0.03 0.72 $\pm$ 0.05 0.70 $\pm$ 0.07 0.70 $\pm$ 0.05 0.72 $\pm$ 0.05
molhiv 0.32 $\pm$ 0.03 0.46 $\pm$ 0.14 0.37 $\pm$ 0.07 0.62 $\pm$ 0.11 0.57 $\pm$ 0.13
Best 14.3% 0.0% 0.0% 28.6% 57.1%
Positive 57.1% 42.9% 85.7% 85.7%
4.3 Comparison to Other Methods
The selection of valid benchmarks from other works is limited. We start by splitting existing works into two broad categories, generalists that function on any graph domain, and specialists that are designed for a specific domain of graph data. The selection of specialist graph models in chemistry is broad, but not necessarily a fair comparison, as these works do not share the same fundamental aims as ToP. They are something of a best-case for pre-training: abundant pre-training data, without missing features. In the following sections we include specialist GNNs, all pre-trained with constrastive learning, as well as specialist LLMs designed for molecular tasks. We don’t expect ToP models to out-perform these specialists. Instead, we include these results to evaluate how our multi-domain pre-training framework compares to the current state of the art within a domain.
The selection of generalist models is far more limited. GCC is, to the best of the author’s knowledge, the only other generalist GNN work that does not rely on some level of feature transformation using LLMs. This carries the significant caveat that the original work completely eliminates features, and the information they carry, from downstream transfer. We extend the original GCC work somewhat, swapping the input and output heads as with our own models, in order to provide fair comparison against ToP.
The use of LLMs to unify the feature input space is we argue, a level of specialism in-and-of itself, as these datasets cannot be immediately applied to new domains, instead requiring hand-design of a textual schema. LLMs themselves have been argued to be generalists [28], in that any graph structure can function as input given some encoding schema.
4.3.1 Molecular Tasks
In Table 3 we compare ToP to other pre-training schemes across molecular tasks. Results are sourced from the original works. These methods are divided into specialist (domain-specific) and generalist (domain-agnostic) approaches. For domain-specific approaches we include InfoGraph [51], GraphCL [17] and AD-GCL [43], contrastive learning approaches. These models are pre-trained with all features included, and hyper-parameters optimised, over Zinc, a dataset of around 2M molecules. We also include MolXPT [27], InstructMol [4] and ChemDFM [58], LLM-based approaches.
Table 3: Our results for the ToP-All and ToP-Social models presented alongside the same results from other pre-training works. Alongside dataset names we include the total size of the dataset. Results are ROC-AUC, and other than for ToP and GCC, taken verbatim from the respective work. Where standard deviations or errors are not present, they were also not present in the respective work. The table is organised by specialism and type (Specialist GNN, Specialist with LLM, Generalist LLM, Generalist GNN). ($n$B) indicates $x$ billion parameters. Non-LLMs are GINs with 6 layers and in the region of 7 million parameters. Bold text indicates a superior result within a type of model.
clintox (1329) bbbp (1835) bace (1361) hiv (37014)
InfoGraph (2.4M) 0.70 $\pm$ 0.03 0.69 $\pm$ 0.01 0.76 $\pm$ 0.02 0.76 $\pm$ 0.02
[51]
GraphCL (2.4M) 0.76 $\pm$ 0.03 0.70 $\pm$ 0.01 0.75 $\pm$ 0.01 0.78 $\pm$ 0.01
[17]
AD-GCL (2.4M) 0.80 $\pm$ 0.04 0.70 $\pm$ 0.01 0.79 $\pm$ 0.01 0.78 $\pm$ 0.01
[43]
MolXPT (350M) 0.95 $\pm$ 0.00 0.80 $\pm$ 0.01 0.88 $\pm$ 0.01 0.78 $\pm$ 0.00
[27]
InstructMol (7B) 0.70 0.82 0.69
[4]
ChemDFM (13B) 0.90 0.67 0.78 0.74
[58]
GPT-4 ($>$200B [^1] Best-guess speculation, as OpenAI is closed-source) 0.52 0.62 0.63 0.66
[31]
LLaMa-2 (13B) 0.46 0.60 0.26 0.29
[45]
Galactica (30B) $\dagger$ 0.82 0.60 0.73 0.76
[44]
GCC (untrained) (200k) 0.50 $\pm$ 0.03 0.59 $\pm$ 0.01 0.44 $\pm$ 0.04 0.46 $\pm$ 0.01
GCC (200k) 0.40 $\pm$ 0.07 0.52 $\pm$ 0.02 0.50 $\pm$ 0.04 0.52 $\pm$ 0.01
GCC+Feat (untrained) 0.79 $\pm$ 0.04 0.64 $\pm$ 0.02 0.70 $\pm$ 0.05 0.70 $\pm$ 0.03
GCC+Feat 0.50 $\pm$ 0.04 0.61 $\pm$ 0.03 0.70 $\pm$ 0.1 0.68 $\pm$ 0.05
ToP-Social (2.4M) 0.86 $\pm$ 0.02 0.62 $\pm$ 0.03 0.70 $\pm$ 0.05 0.62 $\pm$ 0.11
ToP-All 0.84 $\pm$ 0.05 0.67 $\pm$ 0.02 0.72 $\pm$ 0.05 0.57 $\pm$ 0.13
We also include results from other domain-agnostic graph models. For LLM-based approaches these chemical tasks these results are sourced the ChemDFM work [58]. Performing molecular tasks with these models relies on the same assumptions as the chemistry-specific models, but without fine-tuning specifically on chemical literature [^2] Galactica is trained on scientific literature, including but not limited to chemistry.
Finally we include GCC, to our knowledge the only other framework or approach for producing generalised graph models without relying on LLMs, albeit at the cost of feature inclusion for downstream tasks. We fine-tune the two pre-trained encoders packaged with GCC [^3] Our GCC version, updated to current packages, is available at https://github.com/alexodavies/GCC-Updated. Given that GCC does not provide a method for re-introducing domain features, we take the same approach as with our ToP models, swapping the projection heads for nodes and edges (GCC+Feat). We also include GCC with positional encodings as features, as in the original work (GCC). Alongside the pre-trained GCC models we also include results for an equivalent but un-pre-trained model. Where applicable we use the same hyper-parameters as for our fine-tuning of ToP models. In practise this is a batch size of 512 and a learning rate of 0,001.
We find that AD-GCL performs most strongly among the specialist GNN models, MolXPT strongest among the specialist LLMs, Galactica among the generalist LLMs, and ToP-All among the generalist GNNs. Among specialists MolXPT performs most strongly. Among generalists Galactica and our ToP models perform comparably, though both strongly out-perform other models in both classes.
GCC models show consistent negative transfer, both when using their original positional encoding features and when we apply the ToP process for swapping input heads. On the much larger molhiv dataset the smaller (200k parameters vs 7M) GCC models reach greater performance within our 50 epoch fine-tuning runs. Performance from GCC models is consistently better when we apply techniques from this work to include specialist features in place of positional encodings using the ToP approach. As the GCC encoders are much smaller than out ToP encoders, we suspect that smaller models trained with ToP would reach comparable performance.
4.3.2 Non-Molecular Tasks
We apply two recent LLMs, ChatGPT-4o-mini [31] and Llama-3.2-Instruct (3B) [10], across our validation-only datasets (random, community, trees), as well as the Twitch Egos dataset. We use these smaller versions from their respecitve LLM families to reduce computational cost, which is still drastically higher than our GNN-only approach. From here we denote Llama-3.2-Instruct (3B) as Llama-3. These tasks are chosen as they do not rely on node or edge features, meaning that learning and performance directly follows the use of structural information. We also include GCC with positional encodings as features, as in the original work.
LLM Prompting
Using LLMs for arbitrary graph structures requires defining a schema for natural language representations of graphs. We opt for the ‘incident’ encoding style from Fatemi et al [12]. In this format a graph is encoded in the style nodes: $(1,2,3,...)$ edges: $(1\leftrightarrow 2,2\leftrightarrow 3,...)$. We define prompts for each dataset, aimed to both provide contextual information and make tasks clear, which can be found in Listing 2 in the Appendix. We include further prompting text in an attempt to ensure uniformity across responses: ‘Return a number, with no other text or filler. Do not explain your working. Answer in the format THE ANSWER IS X.’ The prompt structure is then {task description + extra prompt information + graph information}. We show an example in Listing 1, and an example response from Llama-3 in the Appendix in Listing 3.
Listing 1: An example prompt for LLMs on the Trees dataset.
question: This is a Random graph. What is the connection probability between two nodes?
Return a number, with no other text or filler. Do not explain your working. Answer in the format THE ANSWER IS X.
graph: nodes: 0, 1, 2, 3, ...
edges: 0 <-> 23, 0 <-> 17, 1 <-> 5, 1 <-> 16, ...
Table 4: Our results for the ToP-All and ToP-Social models presented alongside LLMs and GCC applied to structure-only datasets. Random, Community and Trees are unseen data for all models. Twitch Egos is AUC-ROC, other datasets are RMSE. Due to cost constraints we do not produce error margins for LLM models.
Twitch Egos Random Community Trees
GPT-4o-mini (8B [^4] Best-guess speculation, as OpenAI is closed-source) 0.50 0.30 0.37 0.28
LLaMa-3.2 (3B) 0.46 0.12 0.10 0.42
GCC (untrained) 0.66 $\pm$ 0.01 0.03 $\pm$ 0.00 0.02 $\pm$ 0.01 0.12 $\pm$ 0.01
GCC (200k) 0.66 $\pm$ 0.01 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00
ToP-Social (7M) 0.74 $\pm$ 0.01 0.04 $\pm$ 0.00 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00
ToP-All 0.79 $\pm$ 0.01 0.04 $\pm$ 0.00 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00
As can be seen in Listing 3, models frequently ignored instructions to return only a number, and instead generated large amounts of text. To extract numbers from generated text, we design a simple process, shown in Appendix in Listing 4.
Findings
LLMs, across all tasks, perform worse than ToP or GCC models. This is despite three of our datasets consisting of basic counting tasks, meaning that they are possible to perform in natural language. Larger LLMs might perform better, but given that these models are already many orders of magnitude larger than ToP or GCC models, we view the LLMs used here as fair benchmarks. On the three synthetic datasets (Random, Community, Trees), with their simple structure-based tasks, ToP models and GCC reach essentially the same performance.
That said, a non-pretrained version of the small GCC encoder produces almost exactly the same performance as the pretrained encoder. This indicates that in this setting GCC pre-training has minimal positive transfer benefits. Also worth noting is that the positional encodings GCC uses as node and edge features are explicitly useful for these tasks. ToP models reach the same performance with no such advantage.
On the real-world Twitch Egos dataset ToP models strongly out-perform the LLMs and GCC. A caveatte is that ToP-Social and ToP-All both have graphs from this dataset in their pre-training data. In Section 4.4.1 we present results for ToP models with the Twitch Ego networks excluded from pre-training data, which under fine-tuning had an ROC-AUC performance of 0.73 $\pm$ 0.03, still significantly better than GCC.
4.3.3 Further Transfer Results
By excluding the output head (again see Figure 2) we are able to transfer ToP pre-trained models onto node and edge level tasks. Here we focus on edge prediction and node classification, and though ToP pre-training does seem to offer some performance increase, the benefits are far smaller than on graph level tasks. These results can be found in the Appendix in Table 12.
Further results can be found in the Appendix in Section C: see Table 8 for transfer with a linear models, Tables (9,10) for random GraphCL-like augmentations, Table 11 for ToP with varied GNN backbones, and Table 12 for node and edge classification.
4.4 Further Experiments
Having demonstrated the efficacy of ToP, we present two more experiments, designed to give some insight into why and how ToP pre-training is effective. The first ensures that our results supporting Hypothesis 2, that out-of-domain data can be more useful than in-domain, are not the result of artefacting. The second investigates the extent to which ToP, during downstream transfer, relies on feature information in comparison to structural information.
4.4.1 Varied Pre-Training Data
While the results we present are statistically significant, there remains the possibility of artefacting. In particular it is feasible that a particular domain in the ToP-Social pre-training data lead to positive transfer, not the overall diversity of data. In order to study and eliminate that possibility we perform a limited ablation over the component dataset within the Social pre-training data. We exclude one of the five pre-training datasets at a time, pre-train an encoder with ToP, then fine-tune ten times across our validation datasets, under the same transfer settings as our main results. These results are presented in Table 5.
Table 5: Scores for fine-tuning models pre-trained on non-molecules, each with one non-molecular pre-training dataset excluded. For the classification datasets we report AUROC and for regression datasets RMSE. Bold text indicates a result performs equally (or better) than the model pre-trained on molecules (Chem). Molecular tasks include domain features in fine-tuning.
Social (Name indicates excluded dataset)
Twitch Neurons Cora Facebook Roads
Facebook 0.29 $\pm$ 0.02 0.14 $\pm$ 0.00 0.84 $\pm$ 0.01 0.14 $\pm$ 0.01 0.22 $\pm$ 0.06
Twitch Egos 0.73 $\pm$ 0.03 0.74 $\pm$ 0.01 0.73 $\pm$ 0.01 0.74 $\pm$ 0.01 0.77 $\pm$ 0.01
Roads 0.19 $\pm$ 0.02 0.11 $\pm$ 0.00 0.12 $\pm$ 0.01 0.11 $\pm$ 0.00 0.18 $\pm$ 0.01
Trees 0.15 $\pm$ 0.01 0.12 $\pm$ 0.00 0.13 $\pm$ 0.01 0.12 $\pm$ 0.00 0.15 $\pm$ 0.01
Community 0.28 $\pm$ 0.03 0.02 $\pm$ 0.00 0.07 $\pm$ 0.01 0.02 $\pm$ 0.01 0.22 $\pm$ 0.03
Random 0.15 $\pm$ 0.01 0.05 $\pm$ 0.00 0.09 $\pm$ 0.02 0.04 $\pm$ 0.00 0.13 $\pm$ 0.02
molfreesolv 3.30 $\pm$ 0.32 3.90 $\pm$ 0.24 3.82 $\pm$ 0.31 3.67 $\pm$ 0.37 3.83 $\pm$ 0.38
molesol 1.03 $\pm$ 0.08 1.03 $\pm$ 0.07 1.26 $\pm$ 0.16 1.13 $\pm$ 0.16 1.17 $\pm$ 0.10
mollipo 0.93 $\pm$ 0.03 0.83 $\pm$ 0.03 0.92 $\pm$ 0.05 0.81 $\pm$ 0.03 0.94 $\pm$ 0.02
molclintox 0.49 $\pm$ 0.03 0.76 $\pm$ 0.13 0.72 $\pm$ 0.10 0.76 $\pm$ 0.18 0.74 $\pm$ 0.06
molbbbp 0.64 $\pm$ 0.02 0.64 $\pm$ 0.02 0.61 $\pm$ 0.03 0.62 $\pm$ 0.02 0.64 $\pm$ 0.03
molbace 0.70 $\pm$ 0.04 0.69 $\pm$ 0.07 0.68 $\pm$ 0.05 0.67 $\pm$ 0.06 0.72 $\pm$ 0.05
molhiv 0.41 $\pm$ 0.06 0.55 $\pm$ 0.11 0.48 $\pm$ 0.13 0.63 $\pm$ 0.07 0.52 $\pm$ 0.11
We find that while performance does in-general drop compared to ToP-Social, it is still consistently higher than ToP-Chem. Some datasets do appear to have a larger benefit for transfer than others - for example, when the Twitch Ego networks are excluded, performance on the molclintox dataset drops to near-random. It is worth noting that we do not scale-up the Social pre-training data when a dataset is included, so these ablation models are all pre-trained on 20% less data than ToP-Chem and ToP-Social.
4.4.2 Features vs. Structure
Assume some imbalance between features and structure, $I(y;X)\neq I(y;E)$. Here we propose a way of investigating this imbalance between feature information and structure information. By degrading the useful information in either (or both of) $X$ and $E$, the degree to which performance relies on one or the other should be apparent. A caveat here is that degrading either will presumably also degrade the information from their interdependence.
Consider some destructive noising process $N_{t}(\cdot)$, that degrades the useful information in either $X$ or $E$. $N_{0}(x)=x$, with no useful information after $N_{T}(\cdot)$.
From here edge noising is denoted $N_{E}=N_{t}(E)$, and node feature noising $N_{X}=N_{t}(X)$ Assume some imperfect model produces predictions $f(X,E)\rightarrow\hat{y}$. Performance is some $h(y,\hat{y})$ between real and predicted values. By sampling $h(y,f(N_{X},N_{E})\rightarrow\hat{y})$, the useful information in both $X$ and $E$, and whether performance requires both to be present, should be observable. We apply this experiment over our molecular benchmarks. Here $X$ includes both node and edge features.
Structure noise
We use random edge removal/addition for structure noise. At $t$, we remove $|E_{r}|=p_{t}\cdot|V|$ edges at random from the graph, $E^{\prime}=E\setminus E_{r}$. We then add an equal number of randomly chosen edges $|E_{a}|=p_{t}\cdot|V|$, for a final noised edgelist $E^{\prime\prime}=E^{\prime}\cup E_{a}$.
This is very close to the procedure taken by graph diffusion models as forward noising processes [46]. At $N_{T}(E)$, the structure is a random graph with the same density as the original.
Feature Noise - Permutation
We employ additive Bernoulli noise across node features. At $t$, we have a probability of permuting the entries across a feature $p_{t}$. Let node (or edge features) across a whole dataset be $X\in\{0,1\}^{N\times D}$ with $N$ points and $D$ features. At $t$ we randomly select $N_{f}=p_{t}\cdot D$ features to permute. Each selected feature is then shuffled, across the whole dataset, removing the useful information from that feature that pertains to each sample, while maintaining the marginal distribution of the given feature.
Results
We fix the test set of each benchmark as un-noised, then perform 8 supervised training runs across 8 increasing noise levels on features and structure. In contrast to our transfer results we perform these supervised training runs only for 25 epochs. When noise is applied to structure, no noise is applied to features, and vice-versa. We present the results in Figure 6 for regression datasets and Figure 10 for classification datasets.
Performance suffers by varying amounts across different datasets. On most datasets the untrained GIN’s response to structural noise is minimal, with little performance change even at complete noise in the graph structure. The untrained GIN’s reponse to feature noise is far larger, in particular on regression datasets (molesol, molfreesolv, mollipo). The Chem model, pre-trained only on molecules, follows the same patterns as the untrained GIN model.
On most datasets the two non-molecular models, Social and All, show large performance drops as both structural and feature noise increases. This is most clear on the datasets where ToP pre-training was most effective, for example molclintox. Interestingly this suggests that ToP pre-training allows models not only to use structural patterns to greater effect, but also integrate information from features more usefully. An additional possibility is that ToP pre-training protects against overfitting to patterns in graph features, hence those features being used to greater effect.
Figure 6: Performance variation for each molecular regression benchmark dataset with increasing noise on structure and features. For each dataset, we benchmark performance for an untrained GIN and our three models pre-trained with ToP (Chem-100, Social-100, All-100).
Figure 10: Performance variation for each molecular classification benchmark dataset with increasing noise on structure and features. For each dataset, we benchmark performance for an untrained GIN and our three models pre-trained with ToP (Chem-100, Social-100, All-100).
5 Discussion
Here we discuss our findings in terms of our hypotheses. This includes discussion on the performance of, and reasonable use cases for, models trained with ToP. Finally we detail explicitly assumptions we make and potential risks to the validity of this work.
5.1 Findings
Pre-training on multiple domains of topologies offers significant performance improvements on downstream tasks (Hypothesis 1). With features excluded we find entirely positive transfer across a range of domains and graph-level tasks using adversarially pre-trained models with ToP. With features included, on popular molecular benchmarks, there are significant performance increases on graph-level tasks compared to fully supervised models. This means, broadly, that ToP can be used without risk of negative transfer.
Further, exposure to a larger variety out-of-domain samples can benefit performance. In Table 1, assuming that the ToP-Chem model has the least varied training data, then the ToP-Social model, then the ToP-All model, performance increases along the same order (Hypothesis 2). Performance between the ToP-Social and ToP-All models is highly comparable, despite the ToP-All model having double the training samples. We expect, though we leave this as an area for future research, that a yet greater breadth of domains in the pre-training data would bring greater performance benefits.
5.2 Intuitions and Comparisons
The most intuitive assumption would be that exposing a molecular representation learner to non-molecular graphs would hinder its performance. Instead the ToP-Social and ToP-All models out-perform the ToP-Chem model on almost all molecular datasets under fine-tuning. This holds true for varied pre-training regimes and model backbones (see Tables 910 and 11 in the Appendix). This shows potential for future work using ToP. Pre-training on graphs could be possible through data scale and diversity in-general, instead of within a specific domain. This has clear parallels to the current trends in generalised pre-training in other fields [30, 47].
Similarly one might expect that including node and edge features during supervised training of a non-pretrained model would likely lead to these domain-specific models out-performing our ToP multi-domain models. Our results instead show that performance is both superior and more consistent from our all-domain model on featured molecular graphs, with performance increasing compared to the same model on non-featured graphs.
In comparison to other generalist graph approaches, ToP models perform strongly, including against science-aimed LLM models many orders of magnitude larger in both training data and cost of use. This allows models, after using ToP, to be transferred to arbitrary tasks on graph structures, without the inherent costs of LLM-based approaches.
We extend, using techniques from this work, GCC models to include domain features. Here pre-trained GCC encoders show consistently negative transfer. This indicates that GCC’s use of positional encodings as node and edge features during pre-training actively inhibits their ability to use the information from domain features, which in most applications of graph models are essential for strong performance. ToP models, in contrast, show large performance increases when domain features are included during downstream fine-tuning compared to when domain features are not included. Positive transfer is at least as common with these features included as when they are not.
Given these findings the reasonable use-cases for models pre-trained with ToP are broad. The potential benefit is two-fold. Firstly, pre-training and fine-tuning requires less data for the downstream task to be viable. This is particularly valuable in areas of data scarcity or expense, as well as offering presumed decreases in training time for the final model. Secondly, as shown by our results in this work, our pre-trained models offer actual performance increase over simple supervised training. Our assumption is that ToP pre-training introduces a strong bias towards structural information, which in many applications is a positive characteristic, in-particular where it inhibits overfitting to patterns in the space of node or edge features. Guarantees on this point are left for future research.
Graph Foundation Models
Liu et al [26] recently proposed definitions for, and expected properties of, graph foundation models. The models produced with ToP satisfy their broad definition, ‘…benefit from the pre-training of broad graph data, and can be adapted to a wide range of downstream graph tasks’. The two further properties proposed by Liu et al [26], emergence and homogenization, are less easy to demonstrate. Liu et al.themselves note that how these properties manifest is an open question, and therefore we similarly leave whether ToP models exhibit these properties as an area for future research.
A useful lens is to compare ToP to the much more mature field of foundation models for natural language, ie LLMs. While the current batch of LLMs are deeply expressive, and capable of performing a massive range of natural language tasks at a high level, earlier iterations of the same fundamental technology were more rudimentary. The best parallel for use-cases of ToP might be BERT [9]. BERT is, by current standards, quite a small LLM, and normally requires fine-tuning to reach good performance on downstream tasks. We’re confident that future graph foundation models will be both more powerful and more flexible in terms of tasks, in particular zero-shot and few-shot learning, and leave whether ToP models are an early version of those GFMs as an open question.
5.3 Limitations
In the full scope of pre-training and downstream tasks for graph data, this work has some limitations. Firstly we use only static, non-hyper graphs, as in many other graph learning works. We do not experiment with novel graph augmentations (or “views”), instead employing only node and edge dropping, both random and adversarial. We show that ToP offers performance benefits with random views, and that adversarial views bolster these performance benefits. While adversarial augmentations do improve pre-training, novel augmentations could be developed that include some techniques present in Hassani and Khasahmadi [17], for example, edge addition alongside edge dropping. Recent works find that such augmentations, including combined additive and subtractive edge-based augmentations, can lead to more robust representation learners [13].
Our method primarily focuses on graph-level tasks, in the same manner as the original AD-GCL work [43]. As such, and again as in that original work, we find that performance benefits on node or edge level tasks are much smaller than on graph-level tasks. For node and edge related tasks, future works can optimise ToP through targeted augmentation schemes. Similarly, we employ a limited set of domain compositions in evaluating ToP. A variety of augmentations, paired with a more detailed study of how different compositions of domain data influence representations, could bolster the expressivity of these topology-pre-trained models.
6 Conclusion
We present ToP, a method for generalised multi-domain pre-training of graph models. ToP relies on excluding node and edge features during pre-training. Features can then be re-included during fine-tuning on downstream tasks.
Through contrastive learning, we use ToP to present graph models capable of positive transfer across domains and tasks. These are an important step towards graph foundation models with use-cases in the manner of models such as BERT [9]. Like BERT, an early foundation language model, ToP models can be fine-tuned with positive transfer on multiple domains and tasks. On 75% of downstream tasks from multiple domains we show significant ($p\leq 0.01$) positive transfer compared to a non-pre-trained model with the same architecture, as expected from our first hypothesis. This is also true when node and edge features are re-introduced during fine-tuning. Compared to a supervised baseline, where results are not significantly better, they are never significantly worse. Additionally we showed that ToP pre-training on a single target domain leads to at-best on-par, and on most tasks significantly worse ($p\leq 0.01$), performance compared to pre-training on multiple non-target domains. This clearly demonstrates that the variety provided by out-of-domain pre-training can in this setting hold more value than in-domain pre-training, as in our second hypothesis.
We anticipate that future work will focus on several key areas: 1) Improved or tailored graph augmentation strategies. 2) Investigation of domain weighting in dataset compositions. 3) A larger pre-training dataset spanning a wider range of domains. 4) Improved performance on node-level and edge-level tasks. 5) Two-stage pre-training with an in-domain, features-included method. 6) Integration and adaption for graph transformers
[…]
Supplementary information
All code for ToP, including data processing, result-reproduction and pre-trained models is available in our GitHub repository https://github.com/alexodavies/general-gcl. Our updated code for GCC, including the inclusion of features in downstream tasks, is available in its own GitHub repository https://github.com/alexodavies/GCC.
[…]
Acknowledgements
AOD and RG acknowledge support from the UK Research and Innovation (UKRI) Centre for Doctoral Training in Interactive Artificial Intelligence Award (EP/S022937/1).
Appendix A Dataset Statistics
Table 6: Statistics for each of our datasets. If used for training, we denote these statistics averaged across the training data. Otherwise we report statistics for their validation (downstream fine-tuning) datasets. All tasks are single-target.
Num. Graphs Number of Nodes Number of Edges Diameter Avg. Clustering Downstream Task
Training Only
molpcba 250000 25.7 $\pm$ 6.28 27.7 $\pm$ 7 13.5 $\pm$ 3.29 0.00128 $\pm$ 0.0115
Cora 50000 59.3 $\pm$ 20.7 119$\pm$ 56.6 10.1 $\pm$ 4.68 0.322 $\pm$ 0.0845
Fly Brain 50000 59.5 $\pm$ 20.8 146$\pm$ 99.6 8.8 $\pm$ 3.49 0.237 $\pm$ 0.0875
Training and Evaluation
Twitch 50000 29.6 $\pm$ 11.1 86.6 $\pm$ 70.7 2$\pm$ 0 0.549 $\pm$ 0.149 Class. Games Played
Facebook 50000 59.5 $\pm$ 20.7 206 $\pm$ 170 10.2 $\pm$ 6.39 0.429 $\pm$ 0.133 Avg. Clustering
Roads 50000 59.5 $\pm$ 20.8 73.9 $\pm$ 28.2 16.4 $\pm$ 5.92 0.0559 $\pm$ 0.0426 Diameter / Num. Nodes
Evaluation Only
molesol 1015 12.7 $\pm$ 6.64 12.9 $\pm$ 7.62 6.81 $\pm$ 3.30 0.000353 $\pm$ 0.00417 Water Solubility
molfreesolv 577 8.46 $\pm$ 3.97 8.02 $\pm$ 4.50 5.02 $\pm$ 2.09 0.0 $\pm$ 0.0 Solvation
mollipo 3780 27.0 $\pm$ 7.44 29.4 $\pm$ 8.22 13.8 $\pm$ 4.03 0.00366 $\pm$ 0.017 Lipophilicity
molclintox 1329 26.3 $\pm$ 15.8 28.0 $\pm$ 17.1 12.4 $\pm$ 6.09 0.00259 $\pm$ 0.0191 Class. Toxicity
molbbbp 1835 23.5 $\pm$ 9.89 25.4 $\pm$ 11.0 11.2 $\pm$ 4.03 0.00284 $\pm$ 0.0278 Class. Penetration
molbace 1361 34.0 $\pm$ 7.89 36.8 $\pm$ 8.13 15.2 $\pm$ 3.14 0.00672 $\pm$ 0.0205 Class. Bioactivity
molhiv 37014 25.3 $\pm$ 12.0 27.3 $\pm$ 13.1 12.0 $\pm$ 5.15 0.00158 $\pm$ 0.0156 Class. Inhibition
Evaluation Only, Unseen Domain
Trees 5000 19.6 $\pm$ 6.96 18.6 $\pm$ 6.96 10.1 $\pm$ 3.11 0$\pm$ 0 Depth
Random 5000 29.3 $\pm$ 10.7 111$\pm$ 82.1 3.88 $\pm$ 1.16 0.227 $\pm$ 0.0998 Density
Community 5000 48$\pm$ 0 323$\pm$ 26.2 3$\pm$ 0.0346 0.408 $\pm$ 0.0256 Inter-Community Density
Appendix B Embedding Space
Figure 11: A UMAP embedding of encodings from an untrained encoder.
Figure 12: A UMAP embedding of encodings from the ToP-Chem model, pre-trained on only molecules.
Figure 13: A UMAP embedding of encodings from the ToP-Social model, pre-trained on only non-molecules.
Figure 14: A UMAP embedding of encodings from the ToP-All model, pre-trained on the whole amalgamated dataset (i.e. molecules and non-molecules).
Table 7: On the combined encodings of the validation sets we compute $R^{2}$ correlation coefficients between the first five principal components (PCA) and common graph-level metrics. High correlations, or the lack thereof, should give some indication of how embedding components represent different graph characteristics. Here we present the most correlated metrics and components for each pre-trained model. Underlined text indicates a strong correlation $|R^{2}|\geq 0.66$, and bold text indicates moderate correlation $0.33\leq|R^{2}|\leq 0.66$. We show up to the first five PCA components, but include only three where explained variance values are minimal.
Variance Ratio Metric Correlations
Most $R^{2}$ Second $R^{2}$ Third $R^{2}$
Untrained
PCA 0 $9.96\times 10^{-01}$ Edges $0.219$ Transitivity $0.072$ Density $0.063$
PCA 1 $3.52\times 10^{-03}$ Edges $-0.555$ Transitivity $-0.276$ Density $-0.243$
PCA 2 $1.13\times 10^{-04}$ Edges $0.39$ Transitivity $0.224$ Density $0.21$
Chem
PCA 0 $9.97\times 10^{-01}$ Edges $0.223$ Transitivity $0.074$ Density $0.065$
PCA 1 $2.60\times 10^{-03}$ Edges $-0.514$ Transitivity $-0.253$ Density $-0.225$
PCA 2 $1.14\times 10^{-05}$ Edges $0.132$ Nodes $0.051$ Transitivity $0.047$
Social
PCA 0 $9.99\times 10^{-01}$ Edges $0.127$ Transitivity $0.032$ Density $0.032$
PCA 1 $9.92\times 10^{-04}$ Nodes 1.00 Edges $0.528$ Density $-0.393$
PCA 2 $4.79\times 10^{-08}$ Edges $0.133$ Density $0.07$ Transitivity $0.054$
All
PCA 0 $7.04\times 10^{-01}$ Nodes 0.999 Edges $0.534$ Density $-0.388$
PCA 1 $1.94\times 10^{-01}$ Edges 0.805 Density 0.737 Diameter $-0.554$
PCA 2 $4.23\times 10^{-02}$ Clustering $0.309$ Transitivity $0.248$ Edges $0.235$
PCA 3 $2.43\times 10^{-02}$ Diameter $0.5$ Clustering $-0.416$ Transitivity $-0.243$
PCA 4 $1.82\times 10^{-02}$ Density $-0.132$ Transitivity $-0.108$ Diameter $-0.088$
       PCA 0        PCA 1        PCA 2
Untrained
Refer to caption Refer to caption Refer to caption
Chem
Refer to caption Refer to caption Refer to caption
Social
Refer to caption Refer to caption Refer to caption
All
Refer to caption Refer to caption Refer to caption
Figure 15: PCA components of the encodings of each model across the whole validation dataset, scattered against the most correlated metrics, as in Table 7.
Appendix C Further Results
C.1 Linear Transfer
Table 8: Scores for linear models applied to the encodings of each validation dataset, with and without node labels. For the classification datasets (ogbg-molclintox and Twitch Egos) we report AUROC, other datasets report RMSE. Some datasets, where results are very similar for each model, are excluded. Bold text indicates a superior result. For clarity the results for a given dataset are scaled by a constant factor.
Untrained Chem Social All Magnitude
molfreesolv 4.21 $\pm$ 4.13 3130 $\pm$ 5990 2.00 $\pm$ 0.869 1.01 $\pm$ 1.02 $10^{3}$
molesol 48.3 $\pm$ 90.2 817 $\pm$ 1150 48.1 $\pm$ 62.9 28.1 $\pm$ 38.1 $10^{0}$
mollipo 29.4 $\pm$ 17.9 14.9 $\pm$ 2.0 4.45 $\pm$ 0.62 5.33 $\pm$ 0.56 $10^{0}$
molclintox 0.487 $\pm$ 0.005 0.440 $\pm$ 0.007 0.500 $\pm$ 0.000 0.500 $\pm$ 0.000 $10^{0}$
Facebook 10.6 $\pm$ 1.7 10.7 $\pm$ 20 7.73 $\pm$ 1.87 4.44 $\pm$ 5 $10^{-3}$
Twitch Egos 0.638 $\pm$ 0.012 0.610 $\pm$ 0.072 0.679 $\pm$ 0.002 0.694 $\pm$ 0.002 $10^{0}$
Roads 263 $\pm$ 0.5 256 $\pm$ 0.0 298 $\pm$ 0.3 2.85 $\pm$ 5 $10^{-5}$
Trees 7.66 $\pm$ 0.16 2.85 $\pm$ 0.0 2.98 $\pm$ 0.03 2.56 $\pm$ 0.05 $10^{-3}$
Community 4.36 $\pm$ 0.05 4.25 $\pm$ 0.00 12.0 $\pm$ 2 4.46 $\pm$ 0.07 $10^{-5}$
Random 6.46 $\pm$ 0.7 3.27 $\pm$ 0.0 1.20 $\pm$ 0.2 3.38 $\pm$ 0.0 $10^{-4}$
Untrained-Labels Chem-Labels Social-Labels All-Labels
molfreesolv 480 $\pm$ 172 $(1.13\pm 0.64)\times 10^{6}$ 90900 $\pm$ 83300 11000 $\pm$ 5600 $10^{0}$
molesol 29.4 $\pm$ 7.8 18.5 $\pm$ 7.3 531 $\pm$ 292 34.1 $\pm$ 33.0 $10^{0}$
mollipo 7.33 $\pm$ 1.37 7.25 $\pm$ 1.01 4.02 $\pm$ 0.92 5.01 $\pm$ 0.53 $10^{0}$
molclintox 0.511 $\pm$ 0.022 0.497 $\pm$ 0.001 0.428 $\pm$ 0.001 0.526 $\pm$ 0.060 $10^{0}$
Facebook 14.6 $\pm$ 11.1 10.8 $\pm$ 1.1 7.82 $\pm$ 1.9 4.43 $\pm$ 6 $10^{-3}$
C.2 Transfer Learning
Table 9: Scores for fine-tuning models trained with random node dropping on each validation dataset, with (Model-labels) and without (Model) node labels. For the classification datasets (ogbg-molclintox and Twitch Egos) we report AUROC, other datasets report RMSE. Some datasets, where results are very similar for each model, are excluded. Bold text indicates a result beats the unsupervised (GIN) baseline. Underlined text indicates a superior result, down to three significant figures. GIN results are duplicated from Table 1.
GIN Chem Social All
molfreesolv 4.978 $\pm$ 0.479 3.838 $\pm$ 0.120 4.041 $\pm$ 0.116 4.157 $\pm$ 0.138
molesol 1.784 $\pm$ 0.094 1.897 $\pm$ 0.075 1.520 $\pm$ 0.066 1.357 $\pm$ 0.037
mollipo 1.143 $\pm$ 0.031 1.123 $\pm$ 0.105 1.015 $\pm$ 0.013 0.983 $\pm$ 0.015
molclintox 0.530 $\pm$ 0.042 0.469 $\pm$ 0.030 0.431 $\pm$ 0.002 0.530 $\pm$ 0.086
molbbbp 0.565 $\pm$ 0.043 0.497 $\pm$ 0.027 0.548 $\pm$ 0.037 0.567 $\pm$ 0.032
molbace 0.677 $\pm$ 0.053 0.740 $\pm$ 0.026 0.686 $\pm$ 0.055 0.679 $\pm$ 0.067
molhiv 0.355 $\pm$ 0.059 0.283 $\pm$ 0.001 0.635 $\pm$ 0.046 0.644 $\pm$ 0.039
Facebook 0.311 $\pm$ 0.065 0.261 $\pm$ 0.101 0.159 $\pm$ 0.015 0.145 $\pm$ 0.012
Twitch Egos 0.72 $\pm$ 0.006 0.460 $\pm$ 0.041 0.744 $\pm$ 0.006 0.743 $\pm$ 0.003
Roads 0.46 $\pm$ 0.073 0.228 $\pm$ 0.118 0.150 $\pm$ 0.015 0.132 $\pm$ 0.018
Trees 0.203 $\pm$ 0.012 0.415 $\pm$ 0.250 0.135 $\pm$ 0.015 0.125 $\pm$ 0.009
Community 0.743 $\pm$ 0.106 0.084 $\pm$ 0.083 0.080 $\pm$ 0.018 0.074 $\pm$ 0.021
Random 0.529 $\pm$ 0.228 0.123 $\pm$ 0.075 0.076 $\pm$ 0.010 0.058 $\pm$ 0.005
GIN-Labels Chem-Labels Social-Labels All-Labels
molfreesolv 4.854 $\pm$ 0.513 3.819 $\pm$ 0.136 4.113 $\pm$ 0.070 3.984 $\pm$ 0.162
molesol 1.707 $\pm$ 0.073 1.809 $\pm$ 0.139 1.480 $\pm$ 0.067 1.378 $\pm$ 0.031
mollipo 1.143 $\pm$ 0.031 1.182 $\pm$ 0.128 0.986 $\pm$ 0.017 0.997 $\pm$ 0.011
molclintox 0.497 $\pm$ 0.038 0.474 $\pm$ 0.036 0.430 $\pm$ 0.002 0.474 $\pm$ 0.056
molbbbp 0.561 $\pm$ 0.034 0.491 $\pm$ 0.032 0.543 $\pm$ 0.016 0.565 $\pm$ 0.036
molbace 0.698 $\pm$ 0.050 0.751 $\pm$ 0.024 0.677 $\pm$ 0.030 0.675 $\pm$ 0.088
molhiv 0.374 $\pm$ 0.052 0.286 $\pm$ 0.008 0.615 $\pm$ 0.047 0.628 $\pm$ 0.044
Facebook 0.749 $\pm$ 0.250 0.354 $\pm$ 0.166 0.151 $\pm$ 0.016 0.156 $\pm$ 0.022
Superior Result 19.0% 19.0% 19.0% 61.9%
Beats Baseline 47.6% 76.2% 90.5%
Table 10: Scores for fine-tuning models trained with random edge dropping on each validation dataset, with (Model-labels) and without (Model) node labels. For the classification datasets (ogbg-molclintox and Twitch Egos) we report AUROC, other datasets report RMSE. Some datasets, where results are very similar for each model, are excluded. Bold text indicates a result beats the unsupervised (GIN) baseline. Underlined text indicates a superior result, down to three significant figures. GIN results are duplicated from Table 1.
GIN Chem Social All
molfreesolv 4.978 $\pm$ 0.479 134.800$\pm$ 225.200 4.102 $\pm$ 0.049 4.197 $\pm$ 0.095
molesol 1.784 $\pm$ 0.094 14.050 $\pm$ 23.760 1.312 $\pm$ 0.094 1.571 $\pm$ 0.059
mollipo 1.129 $\pm$ 0.040 1.034 $\pm$ 0.029 0.989 $\pm$ 0.011 1.068 $\pm$ 0.026
molclintox 0.530 $\pm$ 0.042 0.449 $\pm$ 0.013 0.426 $\pm$ 0.006 0.461 $\pm$ 0.029
molbbbp 0.565 $\pm$ 0.043 0.570 $\pm$ 0.040 0.602 $\pm$ 0.025 0.603 $\pm$ 0.029
molbace 0.677 $\pm$ 0.053 0.768 $\pm$ 0.006 0.783 $\pm$ 0.045 0.756 $\pm$ 0.030
molhiv 0.355 $\pm$ 0.059 0.576 $\pm$ 0.152 0.647 $\pm$ 0.066 0.501 $\pm$ 0.135
Facebook 0.311 $\pm$ 0.065 3.70 $\pm$ 6.95 0.180 $\pm$ 0.019 0.237 $\pm$ 0.063
Twitch Egos 0.72 $\pm$ 0.006 0.717 $\pm$ 0.005 0.753 $\pm$ 0.006 0.739 $\pm$ 0.007
Roads 0.46 $\pm$ 0.073 0.156 $\pm$ 0.007 0.167 $\pm$ 0.039 0.252 $\pm$ 0.029
Trees 0.203 $\pm$ 0.012 0.143 $\pm$ 0.009 0.161 $\pm$ 0.019 0.187 $\pm$ 0.020
Community 0.743 $\pm$ 0.106 0.014 $\pm$ 0.000 0.146 $\pm$ 0.024 0.314 $\pm$ 0.069
Random 0.529 $\pm$ 0.228 0.055 $\pm$ 0.012 0.107 $\pm$ 0.019 0.289 $\pm$ 0.098
GIN-Labels Chem-Labels Social-Labels All-Labels
molfreesolv 4.854 $\pm$ 0.513 103.400$\pm$232.200 4.106 $\pm$ 0.021 4.163 $\pm$ 0.069
molesol 1.707 $\pm$ 0.073 8.006 $\pm$ 7.850 1.358 $\pm$ 0.028 1.597 $\pm$ 0.069
mollipo 1.143 $\pm$ 0.031 1.043 $\pm$ 0.026 0.995 $\pm$ 0.014 1.063 $\pm$ 0.038
molclintox 0.497 $\pm$ 0.038 0.446 $\pm$ 0.020 0.422 $\pm$ 0.009 0.475 $\pm$ 0.031
molbbbp 0.561 $\pm$ 0.034 0.535 $\pm$ 0.079 0.612 $\pm$ 0.015 0.589 $\pm$ 0.031
molbace 0.698 $\pm$ 0.050 0.763 $\pm$ 0.018 0.776 $\pm$ 0.032 0.744 $\pm$ 0.034
molhiv 0.374 $\pm$ 0.052 0.632 $\pm$ 0.087 0.652 $\pm$ 0.100 0.479 $\pm$ 0.102
Facebook 0.749 $\pm$ 0.250 2.80 $\pm$ 4.14 0.179 $\pm$ 0.020 0.263 $\pm$ 0.067
Superior Result 9.5% 19.0% 71.4% 4.8%
Beats Baseline 52.4% 90.5% 90.5%
Table 11: Results for fine-tuning for three different layer architectures (Graph Convolution Network, Graph Attention Network, Graph Isomorphism Network) and pre-training data split on molecular benchmark datasets with node and edge features included. On regression datasets (freesolv, esol, lipo) we report MSE, and on classification datasets (bbbp, clintox, bace, hiv) we report 1 - AUROC.
GCN GAT GIN
None Chem Social All None Chem Social All None Chem Social All
Best 2.46 2.38 1.87 2.29 2.55 3.87 3.45 3.34 3.22 3.66 2.98 2.00
freesolv Mean 2.92 3.01 2.62 2.86 3.02 4.02 3.92 4.16 4 4.16 3.61 3.01
$\pm$ Dev. 0.309 0.456 0.363 0.38 0.292 0.113 0.271 0.549 0.434 0.288 0.428 0.594
Best 1.32 0.881 0.896 1.00 1.15 1.40 1.07 1.08 1.48 1.25 0.856 0.807
esol Mean 1.507 0.977 0.965 1.10 1.33 1.69 1.49 1.93 1.72 1.62 1.03 0.897
$\pm$ Dev. 0.095 0.072 0.055 0.094 0.116 0.192 0.262 0.636 0.245 0.306 0.165 0.042
Best 1.02 0.745 0.739 0.779 0.897 0.911 0.979 0.936 1.02 1.05 0.759 0.759
lipo Mean 1.14 0.782 0.765 0.827 1.11 1.03 1.06 1.03 1.09 1.09 0.798 0.788
$\pm$ Dev. 0.066 0.016 0.027 0.042 0.097 0.05 0.093 0.064 0.053 0.056 0.023 0.016
Best 0.303 0.329 0.318 0.327 0.324 0.329 0.322 0.369 0.357 0.347 0.334 0.290
bbbp Mean 0.335 0.375 0.352 0.387 0.352 0.433 0.391 0.416 0.399 0.391 0.374 0.335
$\pm$ Dev. 0.025 0.029 0.02 0.029 0.022 0.058 0.044 0.041 0.03 0.027 0.026 0.021
Best 0.0783 0.150 0.0879 0.0663 0.0471 0.5 0.5 0.489 0.112 0.444 0.106 0.0543
clintox Mean 0.126 0.532 0.305 0.468 0.157 0.5 0.5 0.498 0.351 0.485 0.138 0.164
$\pm$ Dev. 0.036 0.135 0.201 0.134 0.058 0.002 0 0.003 0.124 0.025 0.023 0.054
Best 0.177 0.239 0.195 0.200 0.156 0.238 0.268 0.270 0.210 0.243 0.232 0.203
bace Mean 0.233 0.292 0.28 0.27 0.243 0.378 0.371 0.339 0.247 0.305 0.304 0.276
$\pm$ Dev. 0.036 0.038 0.067 0.048 0.039 0.085 0.08 0.061 0.025 0.073 0.05 0.048
Best 0.466 0.255 0.249 0.245 0.536 0.286 0.308 0.283 0.611 0.533 0.278 0.272
hiv Mean 0.589 0.371 0.345 0.352 0.63 0.514 0.397 0.509 0.679 0.626 0.383 0.434
$\pm$ Dev. 0.074 0.064 0.119 0.072 0.063 0.166 0.082 0.15 0.032 0.066 0.106 0.131
C.3 Node and Edge Transfer
As in graph-level transfer we replace the original input head with a three-layer MLP for both node classification and edge prediction. For node classification, we use a single-layer MLP as an output head. For edge prediction, the embeddings from the encoder are concatenated for the relevant nodes, then passed to a single-layer MLP.
For edge prediction we sample the two-hop ego network of each node from the original graph. For node prediction we take the three-hop ego network, but limit the scale to five neighbours per node. As larger graphs we employ the Facebook and Cora graphs, as well as the Citeseer and PubMed graphs from the same paper. For edge prediction we also include the Amazon Computers and Photo graphs from Shchur et al [41]. For node classification we also employ the GitHub and LastFM Asia graphs from Rozemberczki et al [37]. For edge prediction we take 5% splits for training and testing edges. For node classification we take a 20% split. The results of this transfer can be found in Table 12.
Table 12: Scores for fine-tuning the complete model with node and edge features for node classification and edge prediction. For both node classification and edge prediction we report accuracy. Formatting is the same as Table 1.
Edge Pred. GIN ToP-All (CL) Chem Social All
Amzn. Comp. 0.85 $\pm$ 0.06 0.82 $\pm$ 0.08 0.84 $\pm$ 0.04 0.86 $\pm$ 0.05 0.86 $\pm$ 0.03
Amzn. Photo. 0.86 $\pm$ 0.04 0.86 $\pm$ 0.05 0.89 $\pm$ 0.03 0.88 $\pm$ 0.05 0.91 $\pm$ 0.03
Citeseer 0.81 $\pm$ 0.01 0.78 $\pm$ 0.02 0.75 $\pm$ 0.02 0.75 $\pm$ 0.01 0.80 $\pm$ 0.01
PubMed 0.97 $\pm$ 0.01 0.98 $\pm$ 0.01 0.96 $\pm$ 0.02 0.98 $\pm$ 0.01 0.97 $\pm$ 0.01
Node Class. GIN - Chem Social All
Citeseer 0.75 $\pm$ 0.01 0.71 $\pm$ 0.05 0.72 $\pm$ 0.04 0.72 $\pm$ 0.01 0.74 $\pm$ 0.01
GitHub 0.89 $\pm$ 0.00 0.89 $\pm$ 0.02 0.89 $\pm$ 0.00 0.88 $\pm$ 0.02 0.89 $\pm$ 0.00
LastFM 0.79 $\pm$ 0.00 0.76 $\pm$ 0.02 0.78 $\pm$ 0.01 0.79 $\pm$ 0.01 0.79 $\pm$ 0.00
PubMed 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00
C.4 LLM Usage
Listing 2: Our task prompts for LLM models.
Twitch Egos: This is the ego network of a twitch streamer, with other nodes being other twitch streamers. Do they play one or multiple games? Answer as a probability that they play multiple games.
Random: This is a random graph. What is connection probability between nodes?
Community: This is a community graph. What is the inter-community connection probability?
Trees: This is a tree graph. How deep is it, normalised (divided) by the number of nodes?
Listing 3: The response to the prompt in Listing 1 from Llama-3.2 (3B). The response contains an explanation, despite our prompting specifying not to include any text beyond the answer.
3.5 THE ANSWER IS 3.5.
## Step 1: Calculate the total number of nodes in the graph.
The total number of nodes in the graph is 11.
## Step 2: Calculate the total number of edges in the graph.
The total number of edges in the graph is 10.
## Step 3: Calculate the average depth of the graph by dividing the total number of edges by the total number of nodes. To calculate the average depth, we divide the total number of edges by the total number of nodes.
average_depth = total_edges / total_nodes = 10 / 11 = 0.909090909.
## Step 4: Round the average depth to two decimal places to obtain the final answer.
Rounded_average_depth = round(average_depth, 2) = 0.91.
## Step 5: Since the answer is required in a specific format, we need to convert the rounded average depth to a whole number as the format requires a number without any text or filler. Since the format requires a number, we need to convert the rounded average depth to a whole number. However, the problem statement requires the answer to be a number without any text or filler, which seems to imply a different interpretation of the format. Following the instructions to the letter as requested, we will provide the answer in the required format without further interpretation.
The final answer is: 3.5 THE ANSWER IS 3.5.
Listing 4: Our process for extracting the predictions of LLMs from generated text.
answer = None
for line in response:
# LLMs frequently had a line following the correct format
if "answer" in line.lower():
answer = first number in string
break
if answer is None:
return None
# All tasks have targets 0 <= y <= 1
if answer > 1:
# Assume percentage
answer /= 100
return answer